### v2

# Rectangular representations and $\lambda$-independence of algebraic monodromy groups

Abstract.
Let $\mathfrak{g}$ be a complex semisimple Lie algebra. We define what it means for a finite dimensional representation of $\mathfrak{g}$ to be rectangular and completely classify faithful rectangular representations. As an application, we obtain new $\lambda$-independence results on the algebraic monodromy groups of compatible systems of $\lambda$-adic Galois representations of number fields.
2020 Mathematics Subject Classification:
11F80, 11F70, 11F22, 17B10, 20G05
1. Introduction
1.1. Rectangular representations
1.1.1.
Let $\psi:\mathfrak{g}\to\operatorname{End}(V)$ be a finite dimensional representation of a complex semisimple Lie algebra $\mathfrak{g}$. The following question is of particular interest.
Q1. If $\psi$ is faithful and $\mathfrak{t}$ is a Cartan subalgebra of $\mathfrak{g}$, to what extent is $(\mathfrak{g},\psi)$ determined by the restriction $\psi|_{\mathfrak{t}}$ (i.e., the formal character)?
When $\psi$ is irreducible, this question is completely answered by Larsen-Pink in the group-theoretic perspective [LP90, Theorem 4]. Without assuming the irreducibility of $\psi$, it is proven that the number of $A_{m}$-factors of $\mathfrak{g}$ for $m\in\mathbb{N}\backslash\{1,2,3,4,5,7,8\}$ (resp. the parity of the number of $A_{4}$-factors of $\mathfrak{g}$) is determined by the formal character $\psi|_{\mathfrak{t}}$ [Hui13, Theorems 2.14, 2.17], where $A_{m}:=\mathfrak{sl}(m+1)$. Inspired by a recent work on type $A_{1}$ automorphic Galois representations [HL24], we study a refinement of the above question; see Q2 below.
For $d\in\mathbb{Z}_{\geq 0}$, define $Z_{d}:=\{-d,-d+2,-d+4,\ldots,d-2,d\}\subset\mathbb{Z}$. The size $|Z_{d}|$ is $d+1$. Let $n$ be the rank of $\mathfrak{g}$. The weight lattice $\Lambda_{\mathfrak{g}}$ of $\mathfrak{g}$ (with respect to $\mathfrak{t}$) is a $\mathbb{Z}$-lattice of the $n$-dimensional real vector space $\Lambda_{\mathfrak{g}}\otimes\mathbb{R}$. The formal character $\psi|_{\mathfrak{t}}$ corresponds to a multiset $\Xi$ of weights in $\Lambda_{\mathfrak{g}}\subset\Lambda_{\mathfrak{g}}\otimes\mathbb{R}$. We say that the representation $\psi:\mathfrak{g}\to\operatorname{End}(V)$ is rectangular (Definition 2.6) if
  1. (a)
    every weight in $\Xi$ is of multiplicity one and
  2. (b)
    there exist an $\mathbb{R}$-isomorphism $\iota:\Lambda_{\mathfrak{g}}\otimes\mathbb{R}\to\mathbb{R}^{n}$ and $d_{1},\ldots,d_{n}\in\mathbb{Z}_{\geq 0}$ such that
    $\iota(\Xi)=Z_{d_{1}}\times Z_{d_{2}}\times\cdots\times Z_{d_{n}}.$
Geometrically, (b) means that the weights $\Xi$ can be arranged into a rectangular shape in some coordinate system of $\Lambda_{\mathfrak{g}}\otimes\mathbb{R}$. The set of lengths of the rectangular representation $\psi$ is defined as the multiset
$\mathscr{L}_{\psi}:=\{|Z_{d_{i}}|=d_{i}+1:~1\leq i\leq n\}.$
[…]
Typical examples of rectangular representations are irreducible representations of $A_{1}\times\cdots\times A_{1}$. If $(\mathfrak{g}_{1},\psi_{1})$ and $(\mathfrak{g}_{2},\psi_{2})$ are rectangular with the set of lengths, respectively, equal to $\mathscr{L}_{\psi_{1}}$ and $\mathscr{L}_{\psi_{2}}$, then the external tensor product $(\mathfrak{g}_{1}\times\mathfrak{g}_{2},\psi_{1}\otimes\psi_{2})$ is rectangular with the set of lengths given by the disjoint union $\mathscr{L}_{\psi_{1}}\sqcup\mathscr{L}_{\psi_{2}}$. A rectangular representation $(\mathfrak{g},\psi)$ is said to be indecomposable if it is not equivalent to an external tensor product of two rectangular representations as above. Below is a natural question.
Q2. Can one classify all faithful rectangular representations of complex semisimple Lie algebras?
1.1.2.
Adopt the conventions: $A_{m}=\mathfrak{sl}(m+1)$ ($m\geq 1)$, $B_{m}=\mathfrak{so}(2m+1)$ ($m\geq 1)$, $C_{m}=\mathfrak{sp}(2m)$ ($m\geq 1)$, and $D_{m}=\mathfrak{so}(2m)$ ($m\geq 2)$ for classical complex Lie algebras. All of these are simple, except for $D_{2}=A_{1}\times A_{1}$. Note also the coincidence $A_{1}=B_{1}=C_{1}$, $B_{2}=C_{2}$, and $D_{3}=A_{3}$. We present a complete answer to Q2.
Theorem 1.1.
[…]
  1. (i)
    There exist faithful rectangular representation $(\mathfrak{g}_{1},\psi_{1})$ and faithful indecomposable hypercubic representations $(\mathfrak{g}_{i},\psi_{i})$ for $2\leq i\leq k$ such that
    (1) $(\mathfrak{g},\psi)=\bigotimes_{i=1}^{k}(\mathfrak{g}_{i},\psi_{i})$
    as an external tensor product.
  2. (ii)
    The rectangular representation $\psi_{1}$ admits an external tensor product of indecomposable hypercubic representations
    (2) $(\mathfrak{g}_{1},\psi_{1})=\bigotimes_{j=1}^{s}(\mathfrak{g}_{1,j},\psi_{1,j})$
    such that $\mathfrak{g}_{1}=\prod_{j=1}^{s}\mathfrak{g}_{1,j}$ is some decomposition and each $\psi_{1,j}$ is one of the following.
    1. (a)
      $(A_{1},\mathrm{Sym}^{r}(\mathrm{Std}))$, $r\in\mathbb{N}$.
    2. (b)
      $(A_{1},\mathrm{Sym}^{r_{1}}(\mathrm{Std})\oplus\mathrm{Sym}^{r_{2}}(\mathrm{Std}))$, $r_{1},r_{2}\in\mathbb{Z}_{\geq 0}$ and $|r_{1}-r_{2}|=1$.
    3. (c)
      $(A_{1}\times A_{1},(\mathrm{Std}\otimes\mathbb{1})\oplus(\mathbb{1}\otimes\mathrm{Std}))=(D_{2},\mathrm{Spin})$.
  3. (iii)
    The hypercubic representation $\psi_{i}$ for $2\leq i\leq k$ is one of the following.
    1. (a)
      $(B_{2},\mathrm{Std}\oplus\mathrm{Spin})$.
    2. (b)
      $(B_{m},\mathrm{Spin})$, $m\geq 2$.
    3. (c)
      $(A_{3},\mathrm{Std}\oplus\mathrm{Std}^{\vee})=(D_{3},\mathrm{Spin})$.
    4. (d)
      $(D_{m},\mathrm{Spin})$, $m\geq 4$.
  4. (iv)
    The external tensor product (1) is unique.
  5. (v)
    The external tensor product (2) is unique up to permutation of the factors in the decomposition $\mathfrak{g}_{1}=\prod_{j=1}^{s}\mathfrak{g}_{1,j}$.
Remark 1.2.
Here, $\mathrm{Std}$ (resp. $\mathrm{Spin}$) denotes the standard (resp. spin) representation, and $\mathbb{1}$ denotes the trivial (one-dimensional) representation. For $m\geq 2$, $(B_{m},\mathrm{Spin})$ is irreducible and $(D_{m},\mathrm{Spin})$ is the direct sum of two irreducible (half-spin) representations.
Note that $(A_{1},\mathrm{Std})\otimes(A_{1},\mathrm{Std})$, $(B_{2},\mathrm{Spin})=(C_{2},\mathrm{Std})$, and $(D_{2},\mathrm{Spin})$ are all hypercubic with the same set of lengths {$,2\}$ 2 }. In Theorem 1.1(iii), the hypercubic $(B_{2},\mathrm{Std}\oplus\mathrm{Spin})$ has length 333, while the rest (spin representations) have length 222.
The following corollaries are immediate.
Corollary 1.3.
If $\psi$ is a faithful rectangular representation of a complex semisimple Lie algebra $\mathfrak{g}$, then each simple factor of $\mathfrak{g}$ must be one of $A_{1}$, $A_{3}$, $B_{m}$ for $m\geq 2$, or $D_{m}$ for $m\geq 4$.
Corollary 1.4.
Let $\psi$ be a faithful rectangular representation of a complex semisimple Lie algebra $\mathfrak{g}$ with $\mathscr{L}_{\psi}$ the set of lengths.
  1. (i)
    If $\mathscr{L}_{\psi}$ contains at most one 222 and at most one 333, then $\mathfrak{g}$ has only $A_{1}$-factors.
  2. (ii)
    If every element of $\mathscr{L}_{\psi}$ is even, and 2 appears at most once in $\mathscr{L}_{\psi}$, then the representation decomposes as an external tensor product
    $(\mathfrak{g},\psi)=\bigotimes_{\ell\in\mathscr{L}_{\psi}}(A_{1},\mathrm{Sym}^{\ell-1}(\mathrm{Std}))$
    and it is irreducible.
Corollary 1.5.
If $\psi$ is a rectangular representation of a complex semisimple Lie algebra $\mathfrak{g}$, then $\psi$ is the direct sum of $2^{t}$ irreducible representations for some integer $t$.
Let $\Psi:\mathbf{G}\to\mathrm{GL}(V)$ be a finite dimensional complex representation of a complex reductive group $\mathbf{G}$ and let $\psi:\mathrm{Lie}(\mathbf{G})\to\operatorname{End}(V)$ be the associated Lie algebra representation. We say that $\Psi$ is rectangular (resp. hypercubic) if this is so for the restriction $\psi|_{\mathrm{Lie}(\mathbf{G})^{\operatorname{ss}}}$, where $\mathrm{Lie}(\mathbf{G})^{\operatorname{ss}}$ is the semisimple part of $\mathrm{Lie}(\mathbf{G})$. Then the set of lengths of $\Psi$ is defined to be the set of lengths of $\psi|_{\mathrm{Lie}(\mathbf{G})^{\operatorname{ss}}}$.
The classification results above are applied to deduce new $\lambda$-independence results on the algebraic monodromy groups $\{\mathbf{G}_{\lambda}\}$ of some compatible system $\{\rho_{\lambda}\}$ of Galois representations.
1.2. Application to Galois representations
1.2.1.
Let $K$ be a number field, $\overline{K}$ be an algebraic closure of $K$, and $\operatorname{Gal}_{K}:=\operatorname{Gal}(\overline{K}/K)$ the absolute Galois group of $K$ (equipped with the profinite topology). Let $E$ be another number field and $\Sigma_{E}$ be the set of finite places of $E$. For each finite place $\lambda\in\Sigma_{E}$, denote by $E_{\lambda}$ the $\lambda$-adic completion of $E$ and by $\overline{E}_{\lambda}$ an algebraic closure of $E_{\lambda}$. An $\lambda$-adic representation of $K$ is a continuous group homomorphism
$\rho_{\lambda}:\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(\overline{E}_{\lambda}).$
The algebraic monodromy group of $\rho_{\lambda}$, denoted $\mathbf{G}_{\lambda}$, is a closed subgroup of $\mathrm{GL}_{n,\overline{E}_{\lambda}}$ defined as the Zariski closure of the Galois image $\rho_{\lambda}(\operatorname{Gal}_{K})$ in $\mathrm{GL}_{n,\overline{E}_{\lambda}}$. If $\rho_{\lambda}(\operatorname{Gal}_{K})\subset\mathrm{GL}_{n}(E_{\lambda})$, then $\mathbf{G}_{\lambda}$ becomes an $E_{\lambda}$-subgroup of $\mathrm{GL}_{n,E_{\lambda}}$. If $\rho_{\lambda}$ is semisimple, then $\mathbf{G}_{\lambda}$ is reductive and $\mathbf{G}_{\lambda}^{\operatorname{der}}:=[\mathbf{G}_{\lambda}^{\circ},\mathbf{G}_{\lambda}^{\circ}]$ (the derived group of the identity component $\mathbf{G}_{\lambda}^{\circ}$) is semisimple. Fix an isomorphism $\iota_{\lambda}:\overline{E}_{\lambda}\simeq\mathbb{C}$ for each $\lambda\in\Sigma_{E}$. We define $\mathbf{G}_{\lambda,\mathbb{C}}:=\mathbf{G}_{\lambda}\times_{\iota_{\lambda}}\mathbb{C}$ and $\mathbf{G}_{\lambda,\mathbb{C}}^{\operatorname{der}}:=\mathbf{G}_{\lambda}^{\operatorname{der}}\times_{\iota_{\lambda}}\mathbb{C}$.
Given a compatible system (Definition 3.1) of semisimple $\lambda$-adic Galois representations
(3) $\{\rho_{\lambda}:\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(\overline{E}_{\lambda})\}_{\lambda\in\Sigma_{E}}$
of $K$ defined over $E$, we are interested in the following question on the system of algebraic monodromy groups $\{\mathbf{G}_{\lambda}\}_{\lambda\in\Sigma_{E}}$.
Q3. ($\lambda$-independence of $\mathbf{G}_{\lambda}\subset\mathrm{GL}_{n,\overline{E}_{\lambda}}$)
  1. (i)
    Is the conjugacy class of the complex reductive subgroup $\mathbf{G}_{\lambda,\mathbb{C}}\subset\mathrm{GL}_{n,\mathbb{C}}$ independent of $\lambda\in\Sigma_{E}$?
  2. (ii)
    Suppose $\rho_{\lambda}(\operatorname{Gal}_{K})\subset\mathrm{GL}_{n}(E_{\lambda})$ for each $\lambda\in\Sigma_{E}$. Does there exist a reductive subgroup $\mathbf{G}\subset\mathrm{GL}_{n,E}$ such that $\mathbf{G}\times_{E}E_{\lambda}$ and $\mathbf{G}_{\lambda}$ are conjugate in $\mathrm{GL}_{n,E_{\lambda}}$ for all $\lambda\in\Sigma_{E}$?
[…]
  1. (A)
    The component group $\mathbf{G}_{\lambda}/\mathbf{G}_{\lambda}^{\circ}$ is independent of $\lambda$ [Ser81].
  2. (B)
    The formal character of the reductive subgroup $\mathbf{G}_{\lambda,\mathbb{C}}\subset\mathrm{GL}_{n,\mathbb{C}}$ is independent of $\lambda$ [^1] If $\mathbf{T}_{\lambda,\mathbb{C}}$ is a maximal torus of $\mathbf{G}_{\lambda,\mathbb{C}}$, then the conjugacy class of $\mathbf{T}_{\lambda,\mathbb{C}}$ in $\mathrm{GL}_{n,\mathbb{C}}$ is independent of $\lambda$. [Ser81].
  3. (C)
    The formal bi-character of the reductive subgroup $\mathbf{G}_{\lambda,\mathbb{C}}\subset\mathrm{GL}_{n,\mathbb{C}}$ is independent of $\lambda$ [^2] If $\mathbf{T}_{\lambda,\mathbb{C}}^{\operatorname{ss}}\subset\mathbf{T}_{\lambda,\mathbb{C}}$ is a chain of tori such that $\mathbf{T}_{\lambda,\mathbb{C}}$ is a maximal torus of $\mathbf{G}_{\lambda,\mathbb{C}}$ and $\mathbf{T}_{\lambda,\mathbb{C}}^{\operatorname{ss}}$ is a maximal torus of $\mathbf{G}_{\lambda,\mathbb{C}}^{\operatorname{der}}$, then the conjugacy class of the chain in $\mathrm{GL}_{n,\mathbb{C}}$ is independent of $\lambda$. [Hui13]. In particular, the formal character of the semisimple subgroup $\mathbf{G}_{\lambda,\mathbb{C}}^{\operatorname{der}}\subset\mathrm{GL}_{n,\mathbb{C}}$ is independent of $\lambda$.
[…]
  1. (D)
    If the faithful representation $\mathbf{G}_{\lambda,\mathbb{C}}\hookrightarrow\mathrm{GL}_{n,\mathbb{C}}$ is rectangular for one $\lambda$, this is true for all $\lambda$. Moreover, the set of length of the rectangular $\mathbf{G}_{\lambda,\mathbb{C}}\hookrightarrow\mathrm{GL}_{n,\mathbb{C}}$ is independent of $\lambda$.
By (D), it makes sense to call the semisimple compatible system $\{\rho_{\lambda}\}$ rectangular if the faithful representation $\mathbf{G}_{\lambda,\mathbb{C}}\hookrightarrow\mathrm{GL}_{n,\mathbb{C}}$ is rectangular for some $\lambda$; the set of lengths $\mathscr{L}_{\rho}$ of the rectangular compatible system $\{\rho_{\lambda}\}$ is defined to be the set of lengths of $\mathbf{G}_{\lambda,\mathbb{C}}\hookrightarrow\mathrm{GL}_{n,\mathbb{C}}$.
1.2.2.
[…]
Theorem 1.6.
Let $K$ be a number field and $\{\rho_{\lambda}:\mathrm{Gal}_{K}\to\mathrm{GL}_{n}(\overline{E}_{\lambda})\}_{\lambda\in\Sigma_{E}}$ be a compatible system of semisimple $\lambda$-adic representations of $K$ defined over a number field $E$. Suppose the following conditions are satisfied.
  1. (a)
    $\{\rho_{\lambda}\}$ is rectangular.
  2. (b)
    The set of lengths $\mathscr{L}_{\rho}$ has only even numbers and at most one 222.
Then the algebraic monodromy group $\mathbf{G}_{\lambda,\mathbb{C}}$ is of type $A_{1}$ (i.e., $\mathrm{Lie}(\mathbf{G}_{\lambda,\mathbb{C}}^{\operatorname{der}})=A_{1}\times\cdots\times A_{1}$), the identity component $\mathbf{G}_{\lambda,\mathbb{C}}^{\circ}$ is irreducible on the ambient space and its conjugacy class in $\mathrm{GL}_{n,\mathbb{C}}$ is independent of $\lambda$.
Now assume $K$ is totally real and $\rho_{\lambda}$ is $\mathrm{GL}_{n}(E_{\lambda})$-valued for all $\lambda$. By adopting the techniques in [Hui23a, Hui23b],[HL24] (combining various Galois theoretic and representation theoretic results), we obtain the following $\lambda$-independence results.
Theorem 1.7.
Let $K$ be a totally real field and $\{\rho_{\lambda}:\mathrm{Gal}_{K}\to\mathrm{GL}_{n}(E_{\lambda})\}_{\lambda\in\Sigma_{E}}$ be a compatible system of semisimple $\lambda$-adic representations of $K$ defined over a number field $E$. Suppose the following conditions are satisfied.
  1. (a)
    $\{\rho_{\lambda}\}$ is strictly compatible and regular.
  2. (b)
    $\{\rho_{\lambda}\}$ is connected.
  3. (c)
    $\{\rho_{\lambda}\}$ is rectangular.
  4. (d)
    The set of lengths $\mathscr{L}_{\rho}$ has at most one 222 and at most one 333.
Then the algebraic monodromy group $\mathbf{G}_{\lambda,\mathbb{C}}$ is of type $A_{1}$ and its conjugacy class in $\mathrm{GL}_{n,\mathbb{C}}$ is independent of $\lambda$.
Remark 1.8.
Here are a few remarks about the conditions of the theorem. Let $\pi$ be a regular algebraic, polarized, cuspidal automorphic representation of $\mathrm{GL}_{n}(\mathbb{A}_{K})$, where $K$ is totally real. If the compatible system is associated to $\pi$ (see e.g., [BLGGT14]), then (a) holds. If, for some $\lambda$, the semisimple group $\mathbf{G}_{\lambda,\mathbb{C}}^{\operatorname{der}}$ is of type $A_{1}$ and acts irreducibly on the ambient space, then (c) holds. If (c) holds and the dimension $n$ is not divisible by 444 and 999, then obviously (d) holds. Note that (b),(c),(d) and Corollary 1.4(i) imply that $\mathbf{G}_{\lambda,\mathbb{C}}$ is connected and of type $A_{1}$ for all $\lambda$. Note that the representation $\mathbf{G}_{\lambda,\mathbb{C}}\hookrightarrow\mathrm{GL}_{n,\mathbb{C}}$ can be reducible.
Corollary 1.9.
Let $K$ be a totally real field and $\{\rho_{\lambda}:\mathrm{Gal}_{K}\to\mathrm{GL}_{n}(E_{\lambda})\}_{\lambda\in\Sigma_{E}}$ be a compatible system of semisimple $\lambda$-adic representations of $K$ defined over a number field $E$. Suppose the following conditions are satisfied.
  1. (a)
    $\{\rho_{\lambda}\}$ is strictly compatible and regular.
  2. (b)
    $\{\rho_{\lambda}\}$ is connected.
  3. (c)
    For some $\lambda_{0}$, $\mathbf{G}_{\lambda_{0},\mathbb{C}}^{\operatorname{der}}$ is of type $A_{1}$ and is irreducible on the ambient space.
  4. (d)
    The set of lengths $\mathscr{L}_{\rho}$ has at most one 222 and at most one 333.
Then the following statements hold.
  1. (i)
    The conjugacy class of $\mathbf{G}_{\lambda,\mathbb{C}}$ in $\mathrm{GL}_{n,\mathbb{C}}$ is independent of $\lambda$.
  2. (ii)
    The representation $\rho_{\lambda}$ is absolutely irreducible for all $\lambda$.
  3. (iii)
    The residual representation of $\rho_{\lambda}$ is absolutely irreducible for almost all $\lambda$.
Remark 1.10.
When we replace 1.9(a) by the stronger condition that $\{\rho_{\lambda}\}$ is associated to the automorphic representation $\pi$ in Remark 1.8, this result is obtained previously in [HL24, Theorem 1.1]. Therefore, Corollary 1.9 can be viewed as a partial generalization of [HL24, Theorem 1.1] from automorphic compatible systems to regular strictly compatible systems.
1.3. Organization of the article
We describe the remaining sections of the article.
[…]
2. Classification of rectangular representations
2.1. Root system and Weyl group of $B_{n}$
Definition 2.1.
For $n\in\mathbb{N}$, equip $\mathbb{R}^{n}$ with the standard Euclidean metric and let :=$\mathcal{B}:=\{e_{1},e_{2},...,e_{n}\}$ be the standard orthonormal basis of $\mathbb{R}^{n}$. A subspace of $\mathbb{R}^{n}$ is said to be standard if it is spanned by some elements in $\mathcal{B}$.
The following lemma is obvious.
Lemma 2.2.
A subspace of $\mathbb{R}^{n}$ generated by some standard subspaces is standard.
We describe the root systems $\Phi$ of the classical Lie algebras $B_{n}=\mathfrak{so}(2n+1)$ ($n\geq 1$), $C_{n}=\mathfrak{sp}(2n)$ (for $n\geq 1$), and $D_{n}=\mathfrak{so}(2n)$ (for $n\geq 2$) as subset of the Euclidean space $\mathbb{R}^{n}$ as follows:
(4) $\begin{split}\Phi_{B_{n}}:&=\{\pm e_{i},\pm e_{i}\pm e_{j}\in\mathbb{R}^{n}:~1\leq i<j\leq n\},\\ \Phi_{C_{n}}:&=\{\pm 2e_{i},\pm e_{i}\pm e_{j}\in\mathbb{R}^{n}:~1\leq i<j\leq n\},\\ \Phi_{D_{n}}:&=\{\pm e_{i}\pm e_{j}\in\mathbb{R}^{n}:~1\leq i<j\leq n\}.\end{split}$
We present the figures of $\Phi_{B_{2}}$, $\Phi_{B_{3}}$ and $\Phi_{D_{3}}=\Phi_{A_{3}}$ and establish some useful facts about $\Phi_{B_{n}}$.
$e_{2}$$e_{1}$ $e_{2}$$e_{1}$$e_{3}$
Figure 1. $\Phi_{B_{2}}$ (left) and $\Phi_{B_{3}}$ (right, and $\Phi_{D_{3}}$ consists of long roots in $\Phi_{B_{3}}$)
Proposition 2.3.
The following assertions hold about the root system $\Phi_{B_{n}}$ of $B_{n}$.
  1. (i)
    Any root $\alpha\in\Phi_{B_{n}}$ in (4) has at most two non-zero coordinates.
  2. (ii)
    If $\alpha\in\Phi_{B_{n}}$, then $-\alpha\in\Phi_{B_{n}}$ and the line spanned by $\alpha$ has exactly two roots.
  3. (iii)
    Suppose $n\geq 2$, and let $\alpha=\epsilon_{1}e_{i}+\epsilon_{2}e_{j}$ and $\beta=\epsilon_{3}e_{k}+\epsilon_{4}e_{l}$ be two linearly independent long roots in $\Phi_{B_{n}}$ (where $\epsilon_{1},\epsilon_{2},\epsilon_{3},\epsilon_{4}\in\{\pm 1\}$), $P$ be the plane spanned by $\{\alpha,\beta\}$, and $d$ be the dimension of the space spanned by $\{e_{i},e_{j},e_{k},e_{l}\}$. If $d=4$ (resp. $d=3$), then $P$ has exactly 444 (resp. 666) roots and all of them are long.
  4. (iv)
    If $n\geq 2$ and $P$ is a plane in $\mathbb{R}^{n}$ containing 888 roots in $\Phi_{B_{n}}$, then $P$ is standard and has exactly 888 roots (444 long and 444 short).
  5. (v)
    If $n\geq 3$ and $Q$ is a 333-space in $\mathbb{R}^{n}$ containing 121212 long roots in $\Phi_{B_{n}}$, then $Q$ is either standard or there exist a four dimensional standard subspace $U$ with basis $\{e_{i},e_{j},e_{l},e_{s}\}$ and $\delta_{1},\delta_{2},\delta_{3}\in\{\pm 1\}$ such that $Q$ is the orthogonal complement of $v:=e_{i}+\delta_{1}e_{j}+\delta_{2}e_{l}+\delta_{3}e_{s}$ in $U$.
Proof.
The assertions (i),(ii),(iii) follow easily from (4).
For (iv), we pick two roots $\alpha,\beta$ that span $P$. If $P$ is not standard, then the pair has either one long root or two long roots. In the first case, we may assume $\alpha=e_{i}$, $\beta=e_{j}+\epsilon e_{k}$, $\epsilon\in\{\pm 1\}$ by (ii), and that $\{e_{i},e_{j},e_{k}\}$ spans a 333-space. Since $P$ contains 888 roots, there exists a root $\gamma\in P$ not belonging to {±$\{\pm\alpha,\pm\beta\}$. Hence, $\gamma=a\alpha+b\beta$ for some non-zero $a,b\in\mathbb{R}$ but this contradicts (i). The second case is impossible by (iii) and the assumption that $P$ contains 888 roots. Therefore, we conclude that $P$ is standard and has exactly 888 roots.
For (v), we pick three long roots $\alpha=e_{i}+\epsilon_{1}e_{j},\beta=e_{k}+\epsilon_{2}e_{l},\gamma=e_{r}+\epsilon_{3}e_{s}$ that span $Q$ by (ii), where $\epsilon_{1},\epsilon_{2},\epsilon_{3}\in\{\pm 1\}$. Let $d$ be the dimension of the space spanned by $\{e_{i},e_{j},e_{k},e_{l},e_{r},e_{s}\}$ and $P_{\alpha,\beta}$, $P_{\alpha,\gamma}$, $P_{\beta,\gamma}$ be the planes spanned by respectively $\{\alpha,\beta\}$, $\{\alpha,\gamma\}$, and $\{\beta,\gamma\}$. If $d=5$ (resp. 666), it follows from linear independence of $\{\alpha,\beta,\gamma\}$ and (i) that every root in $Q$ belongs to $P_{\alpha,\beta}\cup P_{\alpha,\gamma}\cup P_{\beta,\gamma}$, and then (iii) implies that $Q$ has exactly 888 (resp. 666) roots. But this is absurd. If $d=3$, then $Q$ is standard. Now suppose $d=4$. If $P_{\alpha,\beta}$ is standard, then $\{i,j\}=\{k,l\}$, $\{i,j\}\cap\{r,s\}=\emptyset$, and (iii) implies that $Q$ has exactly 666 long roots, which is absurd. Hence, all planes $P_{\alpha,\beta}$, $P_{\alpha,\gamma}$, $P_{\beta,\gamma}$ are not standard. Without loss of generality, we assume $\{i,j\}\cap\{k,l\}=\{k\}$ and $\{i,j\}\cap\{r,s\}=\{r\}$. Then $Q$ is the orthogonal complement of $U$ (spanned by $\{e_{i},e_{j},e_{l},e_{s}\}$) of some $v=e_{i}+\delta_{1}e_{j}+\delta_{2}e_{l}+\delta_{3}e_{s}$. ∎
Remark 2.4.
The orthogonal complement of $v=e_{i}+\delta_{1}e_{j}+\delta_{2}e_{l}+\delta_{3}e_{s}$ in $U$ in (v) above contains exactly 121212 (long) roots.
Proposition 2.5.
Let $W_{B_{n}}$ denote the Weyl group of $B_{n}$.
  1. (i)
    Let $\mathbb{H}_{n}:=[-1,1]^{n}\subset\mathbb{R}^{n}$ be the standard $n$-dimensional hypercube. Then $W_{B_{n}}$ is equal to $\{\phi\in O(\mathbb{R}^{n}):~\phi~\text{preserves}~\mathbb{H}_{n}\}$ and satisfies
    $0\to(\mathbb{Z}/2\mathbb{Z})^{n}\to W_{B_{n}}\to S_{n}\to 0,$
    a (split) short exact sequence of groups where $S_{n}$ is the symmetric group of $n$ elements.
  2. (ii)
    Any reflection $\sigma\in W_{B_{n}}$ is induced from a root $\alpha\in B_{n}$, i.e., the 1-1- 1-eigenspace of $\sigma$ is spanned by $\alpha$.
Proof.
[…]
2.2. Rectangular and hypercubic subsets
Definition 2.6.
For $d\in\mathbb{Z}_{\geq 0}$, define $Z_{d}:=\{-d,-d+2,-d+4,...,d-2,d\}\subset\mathbb{Z}$. Let $V_{\mathbb{R}}$ be an $n$-dimensional real vector space.
  1. (1)
    A subset $\Xi\subset V_{\mathbb{R}}$ is said to be rectangular if there is an $\mathbb{R}$-isomorphism $\iota:V_{\mathbb{R}}\to\mathbb{R}^{n}$ such that $\iota(\Xi)=Z_{d_{1}}\times Z_{d_{2}}\times\cdots\times Z_{d_{n}}$ for some $(d_{1},d_{2},...,d_{n})\in\mathbb{Z}_{\geq 0}^{n}$. In this case, we say that $\Xi$ is isomorphic to $Z_{d_{1}}\times Z_{d_{2}}\times\cdots\times Z_{d_{n}}$ and the multiset
    $\mathscr{L}_{\Xi}:=\{\ell_{i}:=|Z_{d_{i}}|=d_{i}+1:~1\leq i\leq n\}$
    is called the set of lengths [^3] This notation is well-defined by Proposition 2.7 below of $\Xi$.
  2. (2)
    If moreover $d_{1}=d_{2}=\cdots=d_{n}$, then $\Xi$ is said to be hypercubic and $\ell=d_{1}+1$ is called the length of $\Xi$.
By definition, we have $|\Xi|=\prod_{i=1}^{n}\ell_{i}$.
Proposition 2.7.
Let $0\leq d_{1}\leq d_{2}\leq\cdots\leq d_{n}$ and $0\leq d_{1}^{\prime}\leq d_{2}^{\prime}\leq\cdots\leq d_{n}^{\prime}$ be integers. Suppose $\phi:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is an isomorphism taking $\Xi=Z_{d_{1}}\times\cdots\times Z_{d_{n}}$ to $\Xi^{\prime}=Z_{d_{1}^{\prime}}\times\cdots\times Z_{d_{n}^{\prime}}$. The following assertions hold.
  1. (i)
    The numbers of zeros in the multisets $\{d_{1},...,d_{n}\}$ and $\{d_{1}^{\prime},...,d_{n}^{\prime}\}$ are equal.
  2. (ii)
    For all $1\leq i\leq n$, one has $d_{i}=d_{i}^{\prime}$.
  3. (iii)
    If $d_{1},d_{1}^{\prime}>0$, then $\phi(e_{i})=\pm e_{j}$ where $d_{i}=d_{j}^{\prime}$.
Proof.
Since the sum of the number of zeros in $\{d_{1},...,d_{n}\}$ (resp. $\{d_{1}^{\prime},...,d_{n}^{\prime}\}$) and the dimension of $\mathrm{Span}\Xi$ (resp. $\mathrm{Span}\Xi^{\prime}$) is equal to $n$, the assertion (i) follows from $\mathrm{Span}\Xi\simeq\mathrm{Span}\Xi^{\prime}$.
For (ii), we may assume $d_{1},d_{1}^{\prime}>0$ by (i). For a subset $S\subset\mathbb{R}^{n}$, define $m(S):=\{\frac{s_{1}+s_{2}}{2}:~s_{1},s_{2}\in S\}$ as the set of mid-points. If $\phi:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is an isomorphism, then $\phi(m(S))=m(\phi(S))$. Define $M_{d}:=m(Z_{d})=\{x\in\mathbb{Z}:~-d\leq x\leq d\}$ for $d\in\mathbb{N}$ and note that
  • $m(\Xi)=m(Z_{d_{1}})\times m(Z_{d_{2}})\times\cdots\times m(Z_{d_{n}})=M_{d_{1}}\times M_{d_{2}}\times\cdots\times M_{d_{n}}$,
  • $\phi(M_{d_{1}}\times M_{d_{2}}\times\cdots\times M_{d_{n}})=M_{d_{1}^{\prime}}\times M_{d_{2}^{\prime}}\times\cdots\times M_{d_{n}^{\prime}}$, and
  • $\pi_{i}(m(\Xi))\subset m(\Xi)$, where $\pi_{i}:\mathbb{R}^{n}\to\mathbb{R}$ denotes the projection to the $i$th coordinate.
Let $0<k_{1}<k_{2}<\cdots<k_{r}$ and $n_{1},n_{2},...,n_{r}$ be positive integers such that
(5) $m(\Xi)=M_{d_{1}}\times M_{d_{2}}\times\cdots\times M_{d_{n}}=M_{k_{1}}^{n_{1}}\times M_{k_{2}}^{n_{2}}\times\cdots\times M_{k_{r}}^{n_{r}}.$
Let $0<\ell_{1}<\ell_{2}<\cdots<\ell_{s}$ and $m_{1},m_{2},...,m_{s}$ be positive integers such that
(6) $m(\Xi^{\prime})=M_{d_{1}^{\prime}}\times M_{d_{2}^{\prime}}\times\cdots\times M_{d_{n}^{\prime}}=M_{\ell_{1}}^{m_{1}}\times M_{\ell_{2}}^{m_{2}}\times\cdots\times M_{\ell_{s}}^{m_{s}}.$
Lemma 2.8.
In equations (5) and (6), $k_{r}=\ell_{s}$ and $n_{r}=m_{s}$. Moreover,
(7) $\phi(0_{\mathbb{R}^{n-n_{r}}}\times M_{k_{r}}^{n_{r}})=0_{\mathbb{R}^{n-m_{s}}}\times M_{\ell_{s}}^{m_{s}}.$
Proof.
Consider a standard basis vector $e_{i}$ in $0_{\mathbb{R}^{n-n_{r}}}\times M_{k_{r}}^{n_{r}}$. Then
(8) $\phi(e_{i}),\phi(k_{r}e_{i})\hskip 7.22743pt\text{are non-zero in}\hskip 7.22743ptM_{\ell_{1}}^{m_{1}}\times M_{\ell_{2}}^{m_{2}}\times\cdots\times M_{\ell_{s}}^{m_{s}}.$
It follows that $k_{r}\leq\ell_{s}$ and thus $k_{r}=\ell_{s}$ by symmetry. Since $\ell_{1},...,\ell_{s-1}<\ell_{s}=k_{r}$, the assertion (8) implies that $\phi(e_{i})\in 0_{\mathbb{R}^{n-m_{s}}}\times M_{\ell_{s}}^{m_{s}}$ and thus $\phi(0_{\mathbb{R}^{n-n_{r}}}\times M_{k_{r}}^{n_{r}})\subset 0_{\mathbb{R}^{n-m_{s}}}\times M_{\ell_{s}}^{m_{s}}$. We obtain (7) by symmetry and $n_{r}=m_{s}$. ∎
Lemma 2.9.
In equations (5) and (6), one has
(9) $\phi(M_{k_{1}}^{n_{1}}\times\cdots\times M_{k_{r-1}}^{n_{r-1}}\times 0_{\mathbb{R}^{n_{r}}})=M_{\ell_{1}}^{m_{1}}\times\cdots\times M_{\ell_{s-1}}^{m_{s-1}}\times 0_{\mathbb{R}^{m_{s}}}.$
Proof.
Suppose $v\in M_{k_{1}}^{n_{1}}\times\cdots\times M_{k_{r-1}}^{n_{r-1}}\times 0_{\mathbb{R}^{n_{r}}}$. If $\phi(v)$ in (6) is not contained in the right-hand side of (9), then (7) implies that there exists $v^{\prime}\in 0_{\mathbb{R}^{n-n_{r}}}\times M_{k_{r}}^{n_{r}}$ such that $\phi(v+v^{\prime})$ lies outside (6), which is absurd. We obtain (9) by symmetry. ∎
Since Lemma 2.8 and 2.9 imply that the two direct products in (5) and (6) are identical and $\phi$ preserves each component, we obtain the assertion (ii) and it suffices to consider $\Xi=\Xi^{\prime}=L_{d}^{n}$ for (iii).
We would like to check that $\phi$ preserves {±$\{\pm e_{i}:~1\leq i\leq n\}\subset M_{d}^{n}$. The case $n=1$ is trivial, so assume $n\geq 2$. Since $de_{i}\in M_{d}^{n}$, we obtain $d\phi(e_{i})=\phi(de_{i})\in M_{d}^{n}\subset\mathbb{Z}^{n}$. Thus, if $\phi(de_{i})=(x_{1},x_{2},...,x_{n})\in\mathbb{Z}^{n}$ then
  • $|x_{\ell}|\in\{0,d\}$ for all $1\leq\ell\leq n$ and
  • not all $x_{\ell}$ is zero.
[…]
Definition 2.10.
Let $0<d_{1}<d_{2}<\cdots<d_{r}$ and $n_{1},n_{2},...,n_{r}$ be positive integers, where $r\in\mathbb{N}$. Let $n:=n_{1}+n_{2}+\cdots+n_{r}$ and define the automorphism groups
$\Omega_{i}:=\{\phi\in\mathrm{GL}_{n_{i}}(\mathbb{R}):~\phi~\text{preserves}~Z_{d_{i}}^{n_{i}}\subset\mathbb{R}^{n_{i}}\}$
for $1\leq i\leq r$ and
$\Omega:=\{\phi\in\mathrm{GL}_{n}(\mathbb{R}):~\phi~\text{preserves}~Z_{d_{1}}^{n_{1}}\times Z_{d_{2}}^{n_{2}}\times\cdots\times Z_{d_{r}}^{n_{r}}\subset\mathbb{R}^{n}\}.$
Corollary 2.11.
Under the standard Euclidean metric on $\mathbb{R}^{n}$, there are natural identifications
  1. (i)
    $\Omega_{i}=W_{B_{n_{i}}}$ for all $i$ and
  2. (ii)
    $\Omega=\Omega_{1}\times\Omega_{2}\times\cdots\times\Omega_{r}$.
In particular, $\Omega$ is orthogonal.
Proof.
Since $\Omega_{i}$ has $2^{n_{i}}n_{i}!$ elements and preserves $\mathbb{H}_{n_{i}}$ the hypercube by Proposition 2.7(iii), the assertion (i) holds by Proposition 2.5(i). The assertion (ii) follows from Proposition 2.7(iii). ∎
2.3. Rectangular and hypercubic representations
Let $\mathfrak{g}$ be a complex semisimple Lie algebra of rank $n$. Fix a Cartan subalgebra $\mathfrak{t}$ of $\mathfrak{g}$ and denote by $\Lambda_{\mathfrak{g}}$ the weight lattice of $\mathfrak{g}$ with respect to $\mathfrak{t}$, regarded as a rank $n$ lattice in $\Lambda_{\mathfrak{g}}\otimes\mathbb{R}\cong\mathbb{R}^{n}$.
Definition 2.12.
A Lie algebra homomorphism $\psi:\mathfrak{g}\to\operatorname{End}(V)$ of the semisimple Lie algebra $\mathfrak{g}$ is determined by its formal character $\psi|_{\mathfrak{t}}$, which corresponds to a multiset $\Xi$ of weights in $\Lambda_{\mathfrak{g}}$. We identity the formal character with the multiset $\Xi$.
  1. (1)
    We say that $\psi$ is rectangular (resp. hypercubic) if each weight of $\Xi$ is of multiplicity one and the subset $\Xi\subset\Lambda_{\mathfrak{g}}\otimes\mathbb{R}$ is rectangular (resp. hypercubic).
  2. (2)
    In this case, we say that the weights in $\Xi$ are rectangular (resp. hypercubic) and we define the set of lengths $\mathscr{L}_{\psi}$ of $\psi$ to be $\mathscr{L}_{\Xi}$ (resp. the length of $\psi$ to be the length of $\Xi$).
  3. (3)
    A rectangular representation $(\mathfrak{g},\psi)$ is said to be decomposable if there exist rectangular representations $(\mathfrak{g}_{1},\psi_{1})$ and $(\mathfrak{g}_{2},\psi_{2})$ such that $\mathfrak{g}=\mathfrak{g}_{1}\times\mathfrak{g}_{2}$ and $\psi\simeq\psi_{1}\otimes\psi_{2}$ as an external tensor product. If such decomposition does not exist, we say that $(\mathfrak{g},\psi)$ is indecomposable.
Remark 2.13.
If $\psi$ above is faithful and rectangular, then the lengths of $\Xi$ are at least 222 because $\Xi$ spans $\Lambda_{\mathfrak{g}}\otimes\mathbb{R}$.
Example 2.14.
Here are some indecomposable hypercubic representations.
  1. (1)
    The one-dimensional trivial representation is hypercubic of length 111.
  2. (2)
    Any irreducible representation of $A_{1}$.
  3. (3)
    The direct sum $V_{1}\oplus V_{2}$ of two irreducible representations of $A_{1}$ whose dimensions differ by 111.
  4. (4)
[…]
    {(±$\{(\pm 1/2,\pm 1/2,...,\pm 1/2)\in\mathbb{R}^{n}\}$
    if the root system is identified with (4). Note that $(B_{n},\mathrm{Spin})$ is irreducible and $(D_{n},\mathrm{Spin})=(D_{n},\mathrm{Spin}^{+}\oplus\mathrm{Spin}^{-})$ is the sum of two irreducible half-spin representations.
  5. (5)
    The direct sum $\mathrm{Std}\oplus\mathrm{Spin}$ representation of $B_{2}$ is hypercubic of length 333 with formal character $\Xi$ equal to
    {(±$\{(\pm 1,0),(0,\pm 1),(0,0)\}\cup\{(\pm 1/2,\pm 1/2)\}\subset\mathbb{R}^{2}$
    if the root system is identified with (4). Note that the restriction of such representation to the subalgebra $A_{1}\times A_{1}\subset B_{2}$ is the external tensor product $(A_{1},\mathrm{Std}\oplus\mathbb{1})\otimes(A_{1},\mathrm{Std}\oplus\mathbb{1})$.
Remark 2.15.
Theorem 1.1 asserts that any faithful rectangular representation of a semisimple Lie algebra is the external tensor product of these indecomposable hypercubic representations in a unique way.
2.4. Reduction to the hypercubic case
Let $I$ be a finite index set. For each $i\in I$, let $\mathfrak{g}_{i}$ be a complex semisimple Lie algebra with corresponding weight lattice $\Lambda_{\mathfrak{g}_{i}}$ (with respect to a Cartan subalgebra $\mathfrak{t}_{i}$) and Weyl group $W_{\mathfrak{g}_{i}}$. The weight lattice $\Lambda_{\mathfrak{g}}$ of the product $\mathfrak{g}:=\prod_{i\in I}\mathfrak{g}_{i}$ (with respect to Cartan subalgebra $\prod_{i\in I}\mathfrak{t}_{i}$) is identified as $\bigoplus_{i\in I}\Lambda_{\mathfrak{g}_{i}}$. The following lemmas are well-known.
Lemma 2.16.
The action of the Weyl group $W_{\mathfrak{g}}$ of $\mathfrak{g}$ on $\Lambda_{\mathfrak{g}}\otimes\mathbb{R}$ is identified as the action of $\prod_{i\in I}W_{\mathfrak{g}_{i}}$ on $\bigoplus_{i\in I}(\Lambda_{\mathfrak{g}_{i}}\otimes\mathbb{R})$. If $\mathfrak{g}_{i}$ is simple, then the action of $W_{\mathfrak{g}_{i}}$ on $\Lambda_{\mathfrak{g}_{i}}\otimes\mathbb{R}$ is irreducible and non-trivial.
Lemma 2.17.
Suppose $\Lambda_{\mathfrak{g}}\otimes\mathbb{R}$ is equipped with a positive definite metric |||||$\cdot$||| | ⋅ | | such that $W_{\mathfrak{g}}$ is orthogonal. The following assertions hold.
  1. (i)
    The direct sum decomposition $\Lambda_{\mathfrak{g}}\otimes\mathbb{R}=\bigoplus_{i\in I}(\Lambda_{\mathfrak{g}_{i}}\otimes\mathbb{R})$ is orthogonal.
  2. (ii)
    If $\mathfrak{g}_{i}$ is a simple, the restriction of |||||$\cdot$||| | ⋅ | | to the subspace $\Lambda_{\mathfrak{g}_{i}}\otimes\mathbb{R}$ is, up to a positive scalar, induced from the Killing form of $\mathfrak{g}_{i}$.
The following general result is crucial to decomposing representations.
Proposition 2.18.
Let $\mathfrak{g}_{1},\mathfrak{g}_{2}$ be semisimple Lie algebras and $\psi:\mathfrak{g}_{1}\times\mathfrak{g}_{2}\to\operatorname{End}(V)$ be a representation such that the formal character $\Xi\subset\Lambda_{\mathfrak{g}_{1}\times\mathfrak{g}_{2}}=\Lambda_{\mathfrak{g}_{1}}\oplus\Lambda_{\mathfrak{g}_{2}}$ is a direct product of multisets:
(10) $\Xi=\Xi_{1}\times\Xi_{2}:=\{(w_{1},w_{2})\in\Lambda_{\mathfrak{g}_{1}}\oplus\Lambda_{\mathfrak{g}_{2}}:~w_{1}\in\Xi_{1},~w_{2}\in\Xi_{2}\},$
where $\Xi_{1}$ (resp. $\Xi_{2}$) is a finite multiset in $\Lambda_{\mathfrak{g}_{1}}$ (resp. $\Lambda_{\mathfrak{g}_{2}}$). Then there exist representations $\psi_{i}:\mathfrak{g}_{i}\to\operatorname{End}(V_{i})$ with formal character $\Xi_{i}$ for $i=1,2$ such that $\psi$ is isomorphic to
$\psi_{1}\otimes\psi_{2}:\mathfrak{g}_{1}\times\mathfrak{g}_{2}\to\operatorname{End}(V_{1}\otimes V_{2}).$
Proof.
Pick a Weyl chamber $C_{i}$ for $\Lambda_{\mathfrak{g}_{i}}\otimes\mathbb{R}$ and we define $w_{i}\in\Lambda_{\mathfrak{g}_{i}}$ to be dominant if $w_{i}\in C_{i}$, $i=1,2$. The product $C_{1}\times C_{2}$ is a Weyl chamber for $\Lambda_{\mathfrak{g}_{i}}\otimes\mathbb{R}$ and we define $(w_{1},w_{2})\in\Lambda_{\mathfrak{g}_{2}\times\mathfrak{g}_{2}}$ to be dominant if $(w_{1},w_{2})\in C_{1}\times C_{2}$. Hence, $(w_{1},w_{2})$ is dominant if and only if $w_{i}$ is dominant for $i=1,2$.
We proceed by induction on the size $k$ of the multiset $\Xi$. It is obvious for $k=1$. In general, let $(w_{1},w_{2})\in\Xi$ be a highest weight, corresponding to an irreducible subrepresentation $U$ of $V$ by the theorem of highest weight. Then $w_{i}$ is a highest weight in $\Xi_{i}$ for $i=1,2$ by (10), and we obtain $U\simeq U_{1}\otimes U_{2}$ where $U_{i}$ is the irreducible representation of $\mathfrak{g}_{i}$ corresponding to $w_{i}$ for $i=1,2$. Let $\Xi_{U_{i}}$ be the formal character of $U_{i}$. If $\Xi_{U_{1}}=\Xi_{1}$ or $\Xi_{U_{2}}=\Xi_{2}$, then the formal character $\Xi^{\prime}$ of $V-U$ is the direct product $\Xi_{1}\times(\Xi_{2}\backslash\Xi_{U_{2}})$ or $(\Xi_{1}\backslash\Xi_{U_{1}})\times\Xi_{2}$ respectively and we are done by the induction hypothesis. Otherwise, we pick a highest weight $w_{2}^{\prime}\in\Xi_{2}\backslash\Xi_{U_{2}}$ so that $(w_{1},w_{2}^{\prime})\in\Xi^{\prime}$ is highest, corresponding to the irreducible subrepresentation $U^{\prime}\simeq U_{1}\otimes U_{2}^{\prime}$ of $V-U$ (where $U_{2}^{\prime}$ is the irreducible representation of $\mathfrak{g}_{2}$ corresponding to $w_{2}^{\prime}$).
If $\Xi_{2}=\Xi_{U_{2}}\cup\Xi_{U_{2}^{\prime}}$ (multiset union), then the formal character $\Xi^{\prime\prime}$ of $V-U-U^{\prime}$ is the direct product $(\Xi_{1}\backslash\Xi_{U_{1}})\times\Xi_{2}$ and we are done by induction hypothesis. Otherwise, we can find a highest weight $w_{2}^{\prime\prime}$ in $\Xi_{2}\backslash(\Xi_{U_{2}}\cup\Xi_{U_{2}^{\prime}})$ corresponding to the irreducible representation $U_{2}^{\prime\prime}$ such that $U_{1}\otimes U_{2}^{\prime\prime}$ is an irreducible subrepresentation of $V-U-U^{\prime}$. Ultimately, we find (by an inductive argument and $\Xi_{2}$ finite) a representation $V_{2}=U_{2}\oplus U_{2}^{\prime}\oplus U_{2}^{\prime\prime}\oplus\cdots$ of $\mathfrak{g}_{2}$ with formal character $\Xi_{2}$ such that the formal character of $V-U_{1}\otimes V_{2}$ is the direct product $(\Xi_{1}\backslash\Xi_{U_{1}})\times\Xi_{2}$. We are done by the induction hypothesis. ∎
The following result enables us to reduce the problem to the hypercubic case.
Proposition 2.19.
Let $\mathfrak{g}$ be a complex semisimple Lie algebra and $(\mathfrak{g},\psi)$ be a faithful rectangular representation with formal character $\Xi\subset\Lambda_{\mathfrak{g}}\otimes\mathbb{R}$ isomorphic to
(11) $Z_{d_{1}}^{n_{1}}\times Z_{d_{2}}^{n_{2}}\times\cdots\times Z_{d_{r}}^{n_{r}}\subset\mathbb{R}^{n},$
where $0<d_{1}<d_{2}<\cdots<d_{r}$ and $n_{1},n_{2},...,n_{r}$ are positive integers, and $n=n_{1}+n_{2}+\cdots+n_{r}$. Then $\mathfrak{g}=\prod_{1\leq i\leq r}\mathfrak{g}_{i}$ is a product of semisimple Lie algebras $\mathfrak{g}_{i}$ and $\psi=\bigotimes_{1\leq i\leq r}\psi_{i}$ is the external tensor product of the faithful hypercubic representations $(\mathfrak{g}_{i},\psi_{i})$ with formal characters $\Xi_{i}\subset\Lambda_{\mathfrak{g}_{i}}\otimes\mathbb{R}$ isomorphic to $Z_{d_{i}}^{n_{i}}\subset\mathbb{R}^{n_{i}}$.
Proof.
Write $\mathfrak{g}=\prod_{j\in J}\mathfrak{q}_{j}$ as a product of simple Lie algebras. Without loss of generality, suppose $\Xi\subset\Lambda_{\mathfrak{g}}\otimes\mathbb{R}$ is equal to (11) and equip $\mathbb{R}^{n}$ as the standard Euclidean metric. Since $W_{\mathfrak{g}}$ preserves the formal character $\Xi$, the Weyl group $W_{\mathfrak{g}}\subset\Omega$ is orthogonal by Corollary 2.11. For each $1\leq i\leq r$, define the $n_{i}$-dimensional subspace
$Y_{i}:=\text{Span}_{\mathbb{R}}\{e_{j}:~n_{1}+\cdots+n_{i-1}<j\leq n_{1}+\cdots+n_{i-1}+n_{i}\}.$
It follows that $Y_{i}$ is a $W_{\mathfrak{g}}$-subrepresentation of $\mathbb{R}^{n}$ (resp. $W_{\mathfrak{q}_{j}}$ for $j\in J$) by Corollary 2.11(ii). As $\mathfrak{q}_{j}$ is simple, $\Lambda_{\mathfrak{q}_{j}}\otimes\mathbb{R}$ is a subspace of a unique $Y_{i}$ by Lemma 2.16. Therefore, $J$ is partitioned into $J_{1}\cup J_{2}\cup\cdots\cup J_{r}$ such that $Y_{i}=\bigoplus_{j\in J_{i}}(\Lambda_{\mathfrak{q}_{j}}\otimes\mathbb{R})$ (by dimension consideration). Let $\mathfrak{g}_{i}=\prod_{j\in J_{i}}\mathfrak{q}_{j}$ for $1\leq i\leq r$. Since $\mathbb{R}^{n}=Y_{1}\oplus Y_{2}\oplus\cdots\oplus Y_{r}$ and $Y_{i}$ is standard subspace for all $i$, we obtain $\Xi_{i}:=Z_{d_{i}}^{n_{i}}\subset\Lambda_{\mathfrak{g}_{i}}\subset\Lambda_{\mathfrak{g}_{i}}\otimes\mathbb{R}=Y_{i}$ for all $i$. Therefore, we are done by applying Proposition 2.18. ∎
2.5. Extracting $A_{1}$-factors and excluding $A_{2}$, $A_{r}$ ($r\geq 4$), and exceptional factors
Let $\psi:\mathfrak{g}\to\operatorname{End}(V)$ be a faithful hypercubic representation of a semisimple Lie algebra $\mathfrak{g}$ and suppose the formal character $\Xi\subset\Lambda_{\mathfrak{g}}\otimes\mathbb{R}$ is equal to $Z_{d}^{n}\subset\mathbb{R}^{n}$ (Euclidean space).
Proposition 2.20.
If $\mathfrak{g}=\mathfrak{g}_{1}\times\mathfrak{g}_{2}$ where $\mathfrak{g}_{1}$ is a product of $A_{1}$-factors and $\mathfrak{g}_{2}$ has no $A_{1}$-factor, then the following assertions hold.
  1. (i)
    $\Lambda_{\mathfrak{g}_{1}}\otimes\mathbb{R}$ and $\Lambda_{\mathfrak{g}_{2}}\otimes\mathbb{R}$ are standard in $\mathbb{R}^{n}$.
  2. (ii)
    There exist faithful hypercubic representations $\psi_{1}:\mathfrak{g}_{1}\to\operatorname{End}(V_{1})$ and $\psi_{2}:\mathfrak{g}_{2}\to\operatorname{End}(V_{2})$ such that $\psi=\psi_{1}\otimes\psi_{2}$.
  3. (iii)
    Moreover, $\mathfrak{g}_{1}=\prod_{1\leq j\leq s}\mathfrak{g}_{1,j}$ and $\psi_{1}=\bigotimes_{1\leq j\leq s}\psi_{1,j}$ is the external tensor product of the faithful hypercubic representations of $\mathfrak{g}_{1,j}$ for $1\leq j\leq s$, where each $\mathfrak{g}_{1,j}$ is either $A_{1}$ or $A_{1}\times A_{1}$.
Proof.
We prove (i),(ii),(iii) simultaneously by induction on $n$. It is obviously true when $n=1$. Suppose $\mathfrak{g}=A_{1}\times\mathfrak{g}^{\prime}$ and let $\sigma$ be the reflection in $W_{A_{1}}\subset W_{\mathfrak{g}}$. Since $W_{\mathfrak{g}}$ preserves $Z_{d}^{n}$ and thus the hypercube $\mathbb{H}_{n}$, it is a subgroup of $W_{B_{n}}$ (the automorphism group of $\mathbb{H}_{n}$) by Proposition 2.5(i). Then Proposition 2.5(ii) asserts that $\sigma$ is induced from a root $\alpha$ of $B_{n}$. We consider two cases.
If $\alpha$ is a short root, then $\Lambda_{A_{1}}\otimes\mathbb{R}$ and thus $\Lambda_{\mathfrak{g}^{\prime}}\otimes\mathbb{R}$ (by Lemma 2.17(i)) are standard subspaces of $\mathbb{R}^{n}$. By Proposition 2.18, $\psi$ is the external tensor product of a representation of $A_{1}$ and a (hypercubic) representation $\psi^{\prime}$ of $\mathfrak{g}^{\prime}$. The result follows from the induction hypothesis on $\psi^{\prime}$.
If $\alpha$ is a long root, then $W_{\mathfrak{g}^{\prime}}$ preserves the fixed part $\mathbb{H}_{n}^{\sigma}$ of the hypercube under $\sigma$ by Lemmas 2.16 and 2.17. Suppose the long root $\alpha=e_{1}+e_{2}$ without loss of generality. Then
(12) $\mathbb{H}_{n}^{\sigma}=\{(x_{1},x_{2})\in\mathbb{R}^{2}:~x_{1}+x_{2}=0,~|x_{1}|,|x_{2}|\leq 1\}\times\mathbb{H}_{n-2}=:L\times\mathbb{H}_{n-2}.$
Since the line segment $L$ has length $2\sqrt{2}$ different from 222 (length of a side of $\mathbb{H}_{n-2}$), $W_{\mathfrak{g}^{\prime}}$ preserves each factor in (12). Thus, $\text{Span}_{\mathbb{R}}L$ is a one-dimensional subrepresentation of $W_{\mathfrak{g}^{\prime}}$. By Lemma 2.17(i) and dimension consideration, we obtain $\mathfrak{g}^{\prime}=A_{1}\times\mathfrak{g}^{\prime\prime}$ and $W_{A_{1}}$ acts non-trivially on the line. It follows that $\Lambda_{A_{1}\times A_{1}}\otimes\mathbb{R}=\text{Span}_{\mathbb{R}}\{e_{1},e_{2}\}$ and $\Lambda_{\mathfrak{g}^{\prime\prime}}\otimes\mathbb{R}=\text{Span}_{\mathbb{R}}\{e_{i}:~i>2\}$ are standard subspaces of $\mathbb{R}^{n}$. By Proposition 2.18, $\psi$ is the external tensor product of a hypercubic representation of $A_{1}\times A_{1}$ and a hypercubic representation $\psi^{\prime\prime}$ of $\mathfrak{g}^{\prime\prime}$. We are done by induction hypothesis on $\psi^{\prime\prime}$. ∎
Given a simple Lie algebra $\mathfrak{g}$, we list some equal-rank subalgebra $\mathfrak{g}^{\prime}\leq\mathfrak{g}$ (using [GOV94, Table 5]):
  1. (1)
    $A_{1}^{r}\leq B_{r}$ for $r\geq 2$.
  2. (2)
    $A_{1}^{r}\leq C_{r}$ for $r\geq 3$.
  3. (3)
    $A_{1}^{2(r-1)}\times A_{3}\leq D_{2r+1}$ for $r\geq 1$.
  4. (4)
    $A_{1}^{2r}\leq D_{2r}$ for $r\geq 2$.
  5. (5)
    $A_{2}\times A_{2}\times A_{2}\leq E_{6}$,
  6. (6)
    $A_{7}\leq E_{7}$,
  7. (7)
    $A_{8}\leq E_{8}$,
  8. (8)
    $A_{2}\times A_{2}\leq F_{4}$,
  9. (9)
    $A_{2}\leq G_{2}$.
Proposition 2.21.
Let $\mathfrak{q}$ be a simple factor of $\mathfrak{g}$. Then $\mathfrak{q}$ cannot be of type $A_{2}$, $A_{r}$ for any $r\geq 4$, and exceptional type.
Proof.
If $\mathfrak{g}^{\prime}$ is a semisimple subalgebra of $\mathfrak{g}$ of equal-rank, then $\psi|_{\mathfrak{g}^{\prime}}$ is also rectangular. Hence, we may assume each simple factor of $\mathfrak{g}$ is of type $A$ by the list above. By Proposition 2.20, we may further assume that $\mathfrak{g}$ has no $A_{1}$-factor. To prove the result, it suffices show that $\mathfrak{g}$ has only $A_{3}$-factors.
Write $\mathfrak{g}=A_{3}^{r}\times\mathfrak{q}_{1}\times\cdots\mathfrak{q}_{k}$ where each $\mathfrak{q}_{i}$ is equal to $A_{n_{i}}$ such that $n_{i}\in\{2\}\cup\mathbb{N}_{\geq 4}$. It follows from Lemma 2.16 and Corollary 2.11 that
$S_{3}^{r}\times\prod_{i=1}^{k}S_{n_{i}+1}\leq S_{4}^{r}\times\prod_{i=1}^{k}S_{n_{i}+1}=W_{A_{3}}^{r}\times\prod_{i=1}^{k}W_{\mathfrak{q}_{i}}=W_{\mathfrak{g}}\subset\Omega=W_{B_{n}}.$
Since the symmetric group $S_{m}$ does not have non-trivial normal 222-subgroup for $m\in\{3\}\cup\mathbb{N}_{\geq 5}$, the homomorphism (via the inclusions above and Proposition 2.5(i))
$S_{3}^{r}\times\prod_{i=1}^{k}S_{n_{i}+1}\hookrightarrow W_{B_{n}}\to S_{n}$
is injective. Since the rank of $\mathfrak{g}$ is $n=3r+n_{1}+\cdots+n_{k}$, we obtain a contradiction by Proposition 2.22. Therefore, $\mathfrak{g}$ has only $A_{3}$-factors. ∎
Proposition 2.22.
Let $i:S_{n_{1}}\times S_{n_{2}}\times\cdots\times S_{n_{k}}\to S_{n}$ be an injective homomorphism such that $n_{i}>1$ for all $1\leq i\leq k$. Then $n_{1}+n_{2}+\cdots+n_{k}\leq n$.
Proof.
For a finite group $G$, let $\mu(G)$ denote the smallest positive integer $m$ such that there exists an injective homomorphism $G\hookrightarrow S_{m}$. Wright [Wri75] proved that
$\mu(S_{n_{1}}\times S_{n_{2}}\times\cdots\times S_{n_{k}})=\sum_{i=1}^{k}\mu(S_{n_{i}})=\sum_{i=1}^{k}n_{i},$
which completes the proof. ∎
2.6. Proof of Theorem 1.1(i)
By Propositions 2.19, 2.20(ii) and 2.21, we may assume that $\psi$ is hypercubic and
$\mathfrak{g}=\mathfrak{q}_{1}\times\mathfrak{q}_{2}\times\cdots\times\mathfrak{q}_{k}$
is a product of simple Lie algebras such that $\mathfrak{q}_{i}$ is of type $B_{m}$ ($m\geq 2$) or $C_{m}$ ($m\geq 3$) or $D_{m}$ ($m\geq 3$ and $D_{3}=A_{3}$). Suppose $\Xi\subset\Lambda_{\mathfrak{g}}\otimes\mathbb{R}$ is equal to $Z_{d}^{n}\subset\mathbb{R}^{n}$. By Proposition 2.18, it suffices to show that each subspace $\Lambda_{\mathfrak{q}_{i}}\otimes\mathbb{R}\subset\mathbb{R}^{n}$ is standard.
Let $\mathfrak{q}$ be a simple factor of $\mathfrak{g}$. Then $W_{\mathfrak{q}}$ is a subgroup of $W_{\mathfrak{g}}\subset\Omega=W_{B_{n}}$ by Lemma 2.16 and Corollary 2.11. For a reflection $\sigma\in W_{\mathfrak{q}}$, denote the line of 1-1- 1-eigenspace of $\sigma$ by $L_{\sigma}\subset\mathbb{R}^{n}$. We record the following observations.
Lemma 2.23.
The following assertions hold for the subspace $\Lambda_{\mathfrak{q}}\otimes\mathbb{R}\subset\mathbb{R}^{n}$.
  1. (i)
    The subspace $\Lambda_{\mathfrak{q}}\otimes\mathbb{R}$ is spanned by $\{L_{\sigma}:~\sigma~\text{is a reflection in }W_{\mathfrak{q}}\}$.
  2. (ii)
    The subspace $\Lambda_{\mathfrak{q}}\otimes\mathbb{R}$ is spanned by some roots in $\Phi_{B_{n}}\subset\mathbb{R}^{n}$ in (4).
Proof.
The first assertion is obvious, and the second follows from the first assertion and Proposition 2.5(ii). ∎
If $\mathfrak{q}$ is of type $B_{m}$ ($m\geq 2$) or $C_{m}$ ($m\geq 3$), then Lemmas 2.17(ii), 2.23(i), and (4) imply that $\Lambda_{\mathfrak{q}}\otimes\mathbb{R}$ is spanned by certain planes $P$ such that each $P$ contains four distinct lines $L_{\sigma}$ (coming from the configuration of $B_{2}$-root system). Since the four lines correspond to eight roots of $B_{n}$ by Proposition 2.5(ii), such $P$ is standard by Proposition 2.3(iv) and thus $\Lambda_{\mathfrak{q}}\otimes\mathbb{R}$ is standard by Lemma 2.1.
[…]
  1. (1)
    $A_{1}^{2(r-1)}\times A_{3}\leq D_{2r+1}$ for $r\geq 1$.
  2. (2)
    $A_{1}^{2r}\leq D_{2r}$ for $r\geq 2$.
Let $\mathfrak{g}^{\prime}$ be the equal-rank subalgebra of $\mathfrak{g}$ obtained by restricting the simple factor $\mathfrak{q}$ of $\mathfrak{g}$ to $\mathfrak{q}^{\prime}$ and letting all other factors remain unchanged. Since the restriction $\psi|_{\mathfrak{g}^{\prime}}$ is hypercubic and all the type $A_{1}$ factors come from $\mathfrak{g}^{\prime}$, Proposition 2.20(i) implies that $\Lambda_{A_{1}^{2(r-1)}}\otimes\mathbb{R}$ if $m=2r+1$ (resp. $\Lambda_{A_{1}^{2r}}\otimes\mathbb{R}$ if $m=2r$) is standard. Since $\Lambda_{\mathfrak{q}}\otimes\mathbb{R}=\Lambda_{\mathfrak{q}^{\prime}}\otimes\mathbb{R}$ in $\mathbb{R}^{n}$, it is standard if we can show that $\Lambda_{A_{3}}\otimes\mathbb{R}$ is standard.
It remains to consider the case $\mathfrak{q}_{1}=D_{3}=A_{3}$. Lemmas 2.17(ii), 2.23(i), and (4) imply that $Q:=\Lambda_{\mathfrak{q}_{1}}\otimes\mathbb{R}$ is a 333-space spanned by six distinct lines $L_{\sigma}$ such that the angle between any two distinct lines is either $90^{\circ}$ or $60^{\circ}/120^{\circ}$. The six lines contain 121212 roots $\alpha$ of $B_{n}$ by Proposition 2.5(ii). If $\alpha\in L_{\sigma}$ is a short root, then the angle between $L_{\sigma}$ and any other of the five lines $L_{\sigma^{\prime}}$ is a multiple of $45^{\circ}$ by (4). It follows that $L_{\sigma}$ is perpendicular to $L_{\sigma^{\prime}}$ as $60^{\circ}/120^{\circ}$ is not a multiple of $45^{\circ}$. Since $Q$ is a 333-space, ten roots of $B_{n}$ lie on the plane spanned by the five lines, which is absurd (Proposition 2.3(iv)). Hence, the twelve roots $\alpha$ are all long roots of $B_{n}$. If $Q$ is not standard, it follows from Proposition 2.3(v) that there exist a standard 444-space $U$ spanned by $\{e_{i},e_{j},e_{l},e_{s}\}$ and $v=e_{i}+\delta_{1}e_{j}+\delta_{2}e_{l}+\delta_{3}e_{s}$ ($\delta_{1},\delta_{2},\delta_{3}\in\{\pm 1\}$) such that $Q\subset U$ is the orthogonal complement of $v$ in $U$. Then Lemma 2.17(i) implies that
$\mathrm{Span}_{\mathbb{R}}\{e_{t}:~1\leq t\leq n,~t\notin\{i,j,l,s\}\}\oplus\mathbb{R}v=(\Lambda_{\mathfrak{q}_{2}}\otimes\mathbb{R})\oplus\cdots\oplus(\Lambda_{\mathfrak{q}_{k}}\otimes\mathbb{R}).$
But this equation is impossible by Lemma 2.23(ii), Proposition 2.3(i), and the fact that $v$ has four non-zero coordinates. We conclude that $Q$ is standard. ∎
2.7. Proof of Theorem 1.1(ii)
By Propositions 2.19 and 2.20(iii), the rectangular representation $\psi_{1}$ of $\mathfrak{g}_{1}$ is the external tensor product $\bigotimes_{j=1}^{s}\psi_{1,j}$ of the hypercubic representations $\psi_{1,j}$ of $\mathfrak{g}_{1,j}$, where $\mathfrak{g}_{1,j}$ is either $A_{1}$ or $A_{1}\times A_{1}$. If $\mathfrak{g}_{1,j}=A_{1}$, then Theorem 1.1(iia,iib) are the only possibilities since formal character $\Xi_{1,j}$ of $\psi_{1,j}$ is rectangular and each weight is of multiplicity one.
Next, suppose $\mathfrak{g}_{1,j}=A_{1}\times A_{1}$ and $\psi_{1,j}$ is not the external tensor product of two rectangular representations (otherwise, we reduce to previous case). Let $\mathbb{Z}^{2}\subset\mathbb{R}^{2}=\Lambda_{A_{1}\times A_{1}}\otimes\mathbb{R}$ be the weight lattice of $A_{1}\times A_{1}$ so that ×$\mathbb{Z}\times\{0\}$ (resp. $\{0\}\times\mathbb{Z}$) corresponds to the weight lattice of the first (resp. second) $A_{1}$. If $d+1$ (at least 222) is the length of the faithful hypercubic $\psi_{1,j}$, then the formal character is equal to
(13) $\{(x,y)\in\mathbb{Z}^{2}:~|x|+|y|\leq d~~\text{and}~~x+y\equiv d~(\text{mod}~2)\}.$
[…]
Figure 2. (13) for $d=4$
[…]
$\bigoplus_{0\leq i\leq d}(A_{1},\mathrm{Sym}^{i}(\mathrm{Std}))\otimes(A_{1},\mathrm{Sym}^{d-i}(\mathrm{Std}))$
is a subrepresentation of $\psi_{1,j}$. We then obtain
$(d+1)^{2}=\sum_{0\leq i\leq d}(d+1)\leq\sum_{0\leq i\leq d}(i+1)(d-i+1)\leq|\Xi_{1,j}|=(d+1)^{2},$
which implies $d+1=(i+1)(d-i+1)$ for all $i$. By putting $i=1$, we obtain $d=1$ and $(A_{1}\times A_{1},\psi_{1,j})=(A_{1}\times A_{1},(\mathrm{Std}\otimes\mathbb{1})\oplus(\mathbb{1}\otimes\mathrm{Std}))$.∎
2.8. Proof of Theorem 1.1(iii)
2.8.1.
By Theorem 1.1(i) and Proposition 2.21, it suffices to classify hypercubic representations $\psi$ of simple Lie algebra $\mathfrak{q}$, where $\mathfrak{q}$ is $A_{3}$, or $B_{n}$ ($n\geq 2$), or $C_{n}$ ($n\geq 3$), or $D_{n}$ ($n\geq 4$). Note that a hypercubic representation $\psi$ is weight multiplicity-free, i.e., every weight has multiplicity one. We present the classification of irreducible weight multiplicity-free representations of simple Lie algebras, as given by Howe.
Theorem 2.24.
[How92, Theorem 4.6.3] The non-trivial irreducible weight multiplicity-free representations of a simple Lie algebra are those on the following lists.
  1. $A_{n}$:
    1. (i)
      the alternating powers $\bigwedge^{m}(\mathbb{C}^{n+1})$ of the standard representation, and
    2. (ii)
      the symmetric powers $\mathrm{Sym}^{m}(\mathbb{C}^{n+1})$ and $\mathrm{Sym}^{m}(\mathbb{C}^{n+1})^{\vee}$ of the standard representation and its dual.
  2. $B_{n}$:
    1. (i)
      the standard representation on $\mathbb{C}^{2n+1}$, and
    2. (ii)
      the spin representation $\mathrm{Spin}$.
  3. $C_{n}$:
    1. (i)
      the standard representation on $\mathbb{C}^{2n}$, and
    2. (ii)
      if $n=2$ or 333, the last fundamental representation, on $\bigwedge_{\mathrm{prim}}^{2}(\mathbb{C}^{4})$ and $\bigwedge_{\mathrm{prim}}^{3}(\mathbb{C}^{6})$ respectively (of dimensions 555 and 141414).
  4. $D_{n}$:
    1. (i)
      the standard representation on $\mathbb{C}^{2n}$, and
    2. (ii)
      the two half-spin representations $\mathrm{Spin}^{+}$ and $\mathrm{Spin}^{-}$.
  5. $E$:
    1. (i)
      the two 272727-dimensional representations of $E_{6}$, and
    2. (ii)
      the 565656-dimensional representation of $E_{7}$.
  6. $G$:
    1. (i)
      the 777-dimensional representation of $G_{2}$.
There are no non-trivial weight multiplicity-free representations for $E_{8}$ and $F_{4}$.
2.8.2.
We begin by classifying the hypercubic representations for $A_{3}$. The weight lattice of $A_{3}$ is generated by the weights of the standard representation $\mathbb{C}^{4}$, which are four vectors $f_{1},f_{2},f_{3},f_{4}\in\mathbb{R}^{3}$ (Euclidean space) satisfying $f_{1}+f_{2}+f_{3}+f_{4}=0$. They can be seen as four vertices of the cube $[-1/2,1/2]^{3}\subset\mathbb{R}^{3}$. (Figure 3).
[…]
Figure 3. $f_{1},f_{2},f_{3},f_{4}$ in $[-1/2,1/2]^{3}$
Lemma 2.25.
Suppose $\Xi=Z_{d_{1}}\times Z_{d_{2}}\times\cdots\times Z_{d_{n}}\subset\mathbb{R}^{n}$ is rectangular and $L_{1},\ldots,L_{r}\subset\mathbb{R}^{n}$ are $r$ affine lines (not necessarily passing through the origin) that satisfy the conditions below.
  1. (a)
    $\min_{1\leq i\leq r}|L_{i}\cap\Xi|=m$,
  2. (b)
    For any $1\leq j\leq n$, there exists $k\in\{1,2,\ldots,n\}$ such that $\pi_{j}(L_{k})=\mathbb{R}$, where $\pi_{j}:\mathbb{R}^{n}\to\mathbb{R}$ denotes projection to the $j$th component.
Then, $|\Xi|\geq m^{n}$.
Proof.
One has from $\pi_{1}(L_{i_{1}}),\ldots,\pi_{n}(L_{i_{n}})=\mathbb{R}$ for some indices $1\leq i_{1},\ldots,i_{n}\leq r$ that
$|Z_{d_{j}}|=|\pi_{j}(L_{i_{j}})\cap\pi_{j}(\Xi)|\geq|\pi_{j}(L_{i_{j}}\cap\Xi)|=|L_{i_{j}}\cap\Xi|\geq m,\quad 1\leq j\leq n,$
and consequently $|\Xi|\geq\prod_{1\leq j\leq n}|Z_{d_{j}}|\geq m^{n}$. ∎
Proposition 2.26.
The only faithful hypercubic representation $V$ of $A_{3}$ is $\mathrm{Std}\oplus\mathrm{Std}^{\vee}$.
Proof.
Let $\Xi\subset\mathbb{R}^{3}$ be the formal character of $V$ that is hypercubic of length $\ell\geq 2$. By Theorem 2.24, if $V$ is hypercubic, then each irreducible factor $W$ of $V$ is of multiplicity one and is equal to $\mathrm{Sym}^{m}(\mathrm{Std})$, or $\mathrm{Sym}^{m}(\mathrm{Std}^{\vee})$, or $\bigwedge^{2}(\mathrm{Std})$, or the trivial representation $\mathbb{1}$. These irreducible representations can be classified as follows: Let $k$ be a non-negative integer.
  1. (1)
    $\mathrm{Sym}^{4k}(\mathrm{Std})$ and $\mathrm{Sym}^{4k}(\mathrm{Std}^{\vee})$ has weight 0, as $0=k(f_{1}+f_{2}+f_{3}+f_{4})$.
  2. (2)
    $\mathrm{Sym}^{4k+1}(\mathrm{Std})$ and $\mathrm{Sym}^{4k+3}(\mathrm{Std}^{\vee})$ has weight $w_{1}:=f_{1}$, as $f_{1}=f_{1}+k(f_{1}+f_{2}+f_{3}+f_{4})=-f_{2}-f_{3}-f_{4}-k(f_{1}+f_{2}+f_{3}+f_{4})$.
  3. (3)
    $\mathrm{Sym}^{4k+2}(\mathrm{Std})$, $\mathrm{Sym}^{4k+2}(\mathrm{Std}^{\vee})$ and $\bigwedge^{2}(\mathrm{Std})$ has weight $w_{2}:=f_{1}+f_{2}$, as $f_{1}+f_{2}=f_{1}+f_{2}+k(f_{1}+f_{2}+f_{3}+f_{4})=-f_{3}-f_{4}-k(f_{1}+f_{2}+f_{3}+f_{4})$.
  4. (4)
    $\mathrm{Sym}^{4k+3}(\mathrm{Std})$ and $\mathrm{Sym}^{4k+1}(\mathrm{Std}^{\vee})$ has weight $w_{3}:=f_{1}+f_{2}+f_{3}$, as $f_{1}+f_{2}+f_{3}=f_{1}+f_{2}+f_{3}+k(f_{1}+f_{2}+f_{3}+f_{4})=-f_{4}-k(f_{1}+f_{2}+f_{3}+f_{4})$.
Since $V$ is hypercubic, $w\in\Xi$ if and only if $-w\in\Xi$ and thus $V$ is self-dual. Hence, if $\mathrm{Sym}^{4k+2}(\mathrm{Std})$ (or $\mathrm{Sym}^{4k+2}(\mathrm{Std}^{\vee})$) is an irreducible factor of $V$ then the weight $w_{2}$ has multiplicity at least 222 in $\Xi$ which is absurd. Similarly, $\mathrm{Sym}^{4k}(\mathrm{Std})$ (or $\mathrm{Sym}^{4k}(\mathrm{Std}^{\vee})$) for $k>0$ cannot be an irreducible factor of $V$. By the facts that $V$ is self-dual and weight multiplicity-free, and the above classification, $V$ is one of the following cases, where $m\in\mathbb{N}$ is congruent to 111 or 333 modulo 444.
  1. (1)
    $V=\mathrm{Sym}^{m}(\mathrm{Std})\oplus\mathrm{Sym}^{m}(\mathrm{Std}^{\vee})\oplus\bigwedge^{2}(\mathrm{Std})\oplus\mathbb{1}$.
  2. (2)
    $V=\mathrm{Sym}^{m}(\mathrm{Std})\oplus\mathrm{Sym}^{m}(\mathrm{Std}^{\vee})\oplus\bigwedge^{2}(\mathrm{Std})$.
  3. (3)
    $V=\mathrm{Sym}^{m}(\mathrm{Std})\oplus\mathrm{Sym}^{m}(\mathrm{Std}^{\vee})\oplus\mathbb{1}$.
  4. (4)
    $V=\mathrm{Sym}^{m}(\mathrm{Std})\oplus\mathrm{Sym}^{m}(\mathrm{Std}^{\vee})$.
  5. (5)
    $V=\bigwedge^{2}(\mathrm{Std})\oplus\mathbb{1}$.
  6. (6)
    $V=\bigwedge^{2}(\mathrm{Std})$.
Among these types of representations, the last two representations $\bigwedge^{2}(\mathrm{Std})\oplus\mathbb{1}$ and $\bigwedge^{2}(\mathrm{Std})$ are 777- and 666-dimensional, respectively, and thus are not hypercubic.
Suppose that $V=\mathrm{Sym}^{m}(\mathrm{Std})\oplus\mathrm{Sym}^{m}(\mathrm{Std}^{\vee})\oplus U$ is one of the first four cases. Let $L_{i}\subset\mathbb{R}^{3}$ be the affine line containing
(14) $\{mf_{4},(m-1)f_{4}+f_{i},\ldots,f_{4}+(m-1)f_{i},mf_{i}\}.$
These are some weights coming from $\mathrm{Sym}^{m}(\mathrm{Std})$ for $i=1,2,3$. Since (14) is a subset of $\Xi$, we obtain
(15) $|L_{i}\cap\Xi|\geq(m+1)\hskip 7.22743pt\text{for}\hskip 7.22743pti=1,2,3.$
Suppose $\Xi=Z_{d_{1}}\times Z_{d_{2}}\times Z_{d_{3}}\subset\mathbb{R}^{3}$ up to a change of coordinates. Since the three lines $L_{1},L_{2},L_{3}$ all pass through $mf_{4}$ and are not coplanar in $\mathbb{R}^{3}$, the condition 2.25(b) is satisfied and we obtain $|\Xi|\geq(m+1)^{3}$ by (15) and the lemma. Hence, we obtain
$\begin{split}\dim V=2\binom{m+3}{3}+\dim U=\frac{1}{3}(m+1)(m+2)(m+3)+\dim U\geq(m+1)^{3},\end{split}$
equivalent to
$\frac{(m+1)(2m+3)(m-1)}{3}\leq\dim U.$
As $\dim U\in\{0,1,6,7\}$ and $m\in\mathbb{N}$ is congruent 111 or 333 modulo 444, the only solution is $m=1$ which implies $\dim V\in\{8,9,14,15\}$. Since $V$ is a faithful hypercubic representation of $A_{3}$, the dimension is given by $\dim V=\ell^{3}$ for some positive integer $\ell\geq 2$ and the only possibility is $8=2^{3}$. In this case, $V=\mathrm{Std}\oplus\mathrm{Std}^{\vee}$ and the formal character is given by
$\Xi=\{\pm f_{1},\pm f_{2},\pm f_{3},\pm f_{4}\}=\{(\pm 1/2,\pm 1/2,\pm 1/2)\}\subset\mathbb{R}^{3},$
which is hypercubic. ∎
2.8.3.
As $D_{3}$ and $A_{3}$ are isomorphic, it remains to classify faithful hypercubic representations $V$ of $B_{n}$ $(n\geq 2)$, $C_{n}$ $(n\geq 3)$, and $D_{n}$ $(n\geq 4)$.
Firstly, consider $B_{n}$ for $n\geq 2$. Since $V$ is weight multiplicity-free, it follows from Theorem 2.24 that
$V=\delta_{1}\mathbb{1}\oplus\delta_{2}\mathrm{Std}\oplus\delta_{3}\mathrm{Spin}$
for $\delta_{1},\delta_{2},\delta_{3}\in\{0,1\}$. If $n>2$, then we have
$\dim(V)\leq\dim(\mathbb{1})+\dim(\mathrm{Std})+\dim(\mathrm{Spin})=2+2n+2^{n}<3^{n},$
so $\dim(V)=2^{n}$. Consequently, $V$ is either $\mathrm{Spin}$, or $n=3$ and $V=\mathrm{Std}\oplus\mathbb{1}$. But the latter one is not hypercubic (and in fact not even weight multiplicity-free), so $V=\mathrm{Spin}$ is the unique hypercubic representation in this case. If $n=2$, then there is an exceptional hypercubic representation $\mathrm{Std}\oplus\mathrm{Spin}$ (see Example 2.14), and one can readily derive that this exceptional representation and the spin representation are the only hypercubic representations of $B_{2}$.
Secondly, consider $C_{n}$ for $n\geq 3$. If $n>3$, then $V$ is either $\mathrm{Std}$ or $\mathrm{Std}\oplus\mathbb{1}$ by weight multiplicity-free of $V$ and Theorem 2.24. But both cases are not hypercubic. When $n=3$, we obtain similarly that
$V=\delta_{1}\mathbb{1}\oplus\delta_{2}\mathrm{Std}\oplus\delta_{3}U$
for $\delta_{1},\delta_{2},\delta_{3}\in\{0,1\}$, where $U=\bigwedge_{\mathrm{prim}}^{3}(\mathbb{C}^{6})$ in Theorem 2.24. As $\dim\mathbb{1}=1$, $\dim\mathrm{Std}=6$, and $\dim U=14$, it is impossible that $\dim V=\ell^{3}$ for any integer $\ell\geq 2$. Therefore, there is no hypercubic representations of $C_{n}$ for $n\geq 3$.
Lastly, consider $D_{n}$ for $n\geq 4$. As above, one obtains $V=\delta_{1}\mathbb{1}\oplus\delta_{2}\mathrm{Std}\oplus\delta_{3}\mathrm{Spin}^{+}\oplus\delta_{4}\mathrm{Spin}^{-}$ for $\delta_{1},\delta_{2},\delta_{3},\delta_{4}\in\{0,1\}$ and thus
$\dim V\leq\dim(\mathbb{1})+\dim(\mathrm{Std})+\dim(\mathrm{Spin}^{+})+\dim(\mathrm{Spin}^{-})=1+2n+2^{n}<3^{n}.$
Therefore, we have $\dim(V)=2^{n}$ and consequently $V=\mathrm{Spin}$ is the unique hypercubic representation of $D_{n}$.
2.9. Proofs of Theorem 1.1(iv),(v)
We handle the uniqueness assertions in Theorem 1.1. By the external tensor product (1), the restriction $\psi|_{\mathfrak{g}_{i}}$ is the direct sum of some copies of $\psi_{i}$. It follows from the classification in Theorem 1.1(iii)(a)–(d) that $\psi_{i}$ is determined by $\psi|_{\mathfrak{g}_{i}}$. Therefore, we obtain Theorem 1.1(iv). For Theorem 1.1(v), suppose
(16) $(\mathfrak{g}_{1},\psi_{1})=\bigotimes_{j^{\prime}=1}^{s^{\prime}}(\mathfrak{g}_{1,j^{\prime}},\psi_{1,j^{\prime}})$
is another external tensor product of indecomposable faithful hypercubic representations, where $\mathfrak{g}_{1}=\prod_{j^{\prime}=1}^{s^{\prime}}\mathfrak{g}_{1,j^{\prime}}$ is some decomposition. We now compare (2) with (16).
Lemma 2.27.
The two sets of factors $F_{1}=\{\mathfrak{g}_{1,j}:~1\leq j\leq s\}$ and $F_{1}^{\prime}=\{\mathfrak{g}_{1,j^{\prime}}:~1\leq j^{\prime}\leq s^{\prime}\}$ of $\mathfrak{g}_{1}$ coincide.
Proof.
Let $V_{1}$ be the ambient space of $\psi_{1}$ and $W_{1,j}$ be the maximal subspace of $V_{1}$ such that $\mathfrak{g}_{1,j}$ acts trivially. Here, $\mathfrak{g}_{1,j}$ is either $A_{1}$ or $A_{1}\times A_{1}$ by Theorem 1.1(ii).
Suppose first $\mathfrak{g}_{1,j}=A_{1}$. By the tensor product decomposition (2) and Theorem 1.1(ii)(a),(b), we obtain $\dim W_{1,j}<(\dim V_{1})/2.$ If $\mathfrak{g}_{1,j}$ is a factor of some $\mathfrak{g}_{1,j^{\prime}}=A_{1}\times A_{1}$, then $\dim W_{1,j}=(\dim V_{1})/2$ by (16) and Theorem 1.1(ii)(c), which is absurd. Hence, the simple factors in $F_{1}$ and $F_{1}^{\prime}$ coincide.
Next, we suppose $\mathfrak{g}_{1,j}=A_{1}\times A_{1}$. In this case, we have $\dim W_{1,j}=0$ by the tensor product decomposition (2) and Theorem 1.1(ii)(c). If $\mathfrak{g}_{1,j}\notin F_{1}^{\prime}$, then there exist two distinct non-simple factors $\mathfrak{g}_{1,j^{\prime}}$ and $\mathfrak{g}_{1,j^{\prime\prime}}$ in $F_{1}^{\prime}$ that intersect $\mathfrak{g}_{1,j}$ non-trivially, since the simple factors in $F_{1}$ and $F_{1}^{\prime}$ coincide. This implies that $\dim W_{1,j}\neq 0$ by (16) and Theorem 1.1(ii)(c), which is absurd. Therefore, we conclude that $F_{1}$ and $F_{1}^{\prime}$ coincide. ∎
Lemma 2.27 induces a bijective correspondence $j\leftrightarrow j^{\prime}$ so that $\mathfrak{g}_{1,j}=\mathfrak{g}_{1,j^{\prime}}$ in $\mathfrak{g}_{1}$. Then Theorem 1.1(ii) implies that
$\text{direct sum of some copies of }\psi_{1,j}=\psi_{1}|_{\mathfrak{g}_{1,j}}=\psi_{1}|_{\mathfrak{g}_{1,j^{\prime}}}=\text{direct sum of some copies of }\psi_{1,j^{\prime}}$
and we obtain $\psi_{1,j}=\psi_{1,j^{\prime}}$ from the classification in Theorem 1.1(ii)(a)–(c).
3. $\lambda$-independence of algebraic monodromy groups
3.1. Notation and terminologies
We collect the notation and terminologies that will be used frequently.
  • $K$ (resp. $E$): a number field.
  • $\Sigma_{K}$ (resp. $\Sigma_{E}$): the set of finite places of $K$ (resp. $E$).
  • $v$ (resp. $\lambda$): an element of $\Sigma_{K}$ (resp. $\Sigma_{E}$).
  • $p$ (resp. $\ell$): the residue characteristic of $v$ (resp. $\lambda$).
  • $S_{\ell}$: the set of places in $\Sigma_{K}$ that divide the rational prime $\ell$.
  • $\overline{K}$: an algebraic closure of $K$.
  • $E_{\lambda}$ and $\mathbb{F}_{\lambda}$: the $\lambda$-adic completion of $E$ and the residue field of $E_{\lambda}$.
  • $\overline{E}_{\lambda}$ (resp. $\overline{\mathbb{F}}_{\lambda}$): an algebraic closure of $E_{\lambda}$ (resp. $\mathbb{F}_{\lambda}$).
  • $\operatorname{Gal}_{K}$: the absolute Galois group $\operatorname{Gal}(\overline{K}/K)$ equipped with the profinite topology.
  • $\rho_{\lambda}:\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(\overline{E}_{\lambda})$: an $\lambda$-adic representation of $K$, i.e., a continuous group homomorphism.
  • $\bar{\rho}_{\lambda}^{\operatorname{ss}}:\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(\overline{\mathbb{F}}_{\lambda})$: the residue representation of $\rho_{\lambda}$, i.e., the semisimplification of the reduction modulo $\lambda$ of $\rho_{\lambda}$.
  • $\mathbf{G}_{\lambda}$ or $\mathbf{G}_{\rho_{\lambda}}$: the algebraic monodromy group of $\rho_{\lambda}$, i.e., the Zariski closure of the Galois image $\rho_{\lambda}(\operatorname{Gal}_{K})$ in $\mathrm{GL}_{n,\overline{E}_{\lambda}}$.
  • $\iota_{\lambda}:\overline{E}_{\lambda}\to\mathbb{C}$: a field isomorphism for each $\lambda\in\Sigma_{E}$.
  • $\mathbf{G}_{\lambda,\mathbb{C}}$ (resp. $\mathbf{G}_{\rho_{\lambda},\mathbb{C}}$): the base change $\mathbf{G}_{\lambda}\times_{\iota_{\lambda}}\mathbb{C}$ (resp. $\mathbf{G}_{\rho_{\lambda}}\times_{\iota_{\lambda}}\mathbb{C}$).
  • $\mathbf{G}^{\circ}$: the identity component of a linear algebraic group $\mathbf{G}$.
  • $\mathbf{G}^{\operatorname{der}}$: the derived group $[\mathbf{G}^{\circ},\mathbf{G}^{\circ}]$ of the identity component of $\mathbf{G}$.
  • $\epsilon_{\ell}$ (resp. $\bar{\epsilon}_{\ell}$): the $\ell$-adic (resp. mod $\ell$) cyclotomic character.
  • A semisimple $\lambda$-adic representation $\sigma_{\lambda}$ is of type $A$ (resp. of type $A_{1}$) if the semisimple part $\mathrm{Lie}(\mathbf{G}_{\sigma_{\lambda},\mathbb{C}})^{\operatorname{ss}}$ has only type $A$ (resp. type $A_{1}$) factors. This includes the case where the semisimple part is zero.
  • $\mathbb{G}_{m}$ and $\mathbb{G}_{a}$: $\mathrm{Spec}(F[x,1/x])$ and $\mathrm{Spec}(F[x])$ for some field $F$.
3.2. Compatible system of Galois representations
[…]
Definition 3.1.
A family of ($n$-dimensional) $\lambda$-adic Galois representations indexed by $\Sigma_{E}$,
$\{\rho_{\lambda}:\operatorname{Gal}_{K}\longrightarrow\mathrm{GL}_{n}(\overline{E}_{\lambda})\}_{\lambda\in\Sigma_{E}},$
[…]
  1. (a)
    For each $\lambda\in\Sigma_{E}$, the representation $\rho_{\lambda}$ is unramified at every $v\in\Sigma_{K}\backslash(S\cup S_{\ell})$.
  2. (b)
    For each $\lambda\in\Sigma_{E}$ and $v\in\Sigma_{K}\backslash(S\cup S_{\ell})$, the characteristic polynomial of $\rho_{\lambda}(\mathrm{Frob}_{v})$ satisfies
    $\det(\rho_{\lambda}(\mathrm{Frob}_{v})-T\cdot\mathrm{id})=\Phi_{v}(T)\in E[T].$
The compatible system $\{\rho_{\lambda}\}$ is said to be semisimple if each $\rho_{\lambda}$ is semisimple.
Definition 3.2.
A semisimple $E$-rational compatible system $\{\rho_{\lambda}:\mathrm{Gal}_{K}\to\mathrm{GL}_{n}(\overline{E}_{\lambda})\}_{\lambda\in\Sigma_{E}}$ is called a weakly compatible system if the following conditions are satisfied.
  1. (a)
    For each $\lambda\in\Sigma_{E}$ and $v\in S_{\ell}$, the local representation $\rho_{\lambda}|_{\operatorname{Gal}_{K_{v}}}$ is de Rham and is further crystalline if $v\not\in S$.
  2. (b)
    For each embedding $\tau:K\hookrightarrow\overline{E}$, there is a multiset of $n$ integers $\mathrm{HT}_{\tau}$ such that $\mathrm{HT}_{i\circ\tau}(\rho_{\lambda})=\mathrm{HT}_{\tau}$ for any $\lambda$ and any $i:\overline{E}\hookrightarrow\overline{E}_{\lambda}$ over $E$.
A weakly compatible system is called a strictly compatible system if an extra condition holds.
  1. (c)
    For $\lambda\in\Sigma_{E}$ and $v\notin S_{\ell}$, the semisimplified Weil-Deligne representation $\iota\mathrm{WD}(\rho_{\lambda}|_{\mathrm{Gal}_{K_{v}}})^{F-ss}$ is independent of $\lambda$ and $\iota:\overline{E}_{\lambda}\stackrel{{\scriptstyle\cong}}{{\rightarrow}}\mathbb{C}$.
Definition 3.3.
An $\lambda$-adic representation $\rho_{\lambda}:\mathrm{Gal}_{K}\to\mathrm{GL}_{n}(\overline{E}_{\lambda})$ is said to be regular if it is unramified almost everywhere, $\rho_{\lambda}$ is de Rham at any $v$ above $\ell$, and $\mathrm{HT}_{\tau}(\rho_{\lambda})$ consists of $n$ distinct numbers for any embedding $\tau:K\hookrightarrow\overline{E}_{\lambda}$.
A weakly compatible system $\{\rho_{\lambda}:\mathrm{Gal}_{K}\to\mathrm{GL}_{n}(\overline{E}_{\lambda})\}_{\lambda}$ is said to be regular if some (and hence all) $\rho_{\lambda}$ is regular.
[…]
3.3. Potential automorphy of Galois representations
Given an $\ell$-adic Galois representation $\rho_{\ell}:\mathrm{Gal}_{K}\to\mathrm{GL}_{n}(\overline{\mathbb{Q}}_{\ell})$, one can ask whether $\rho_{\ell}$ is automorphic (and thus part of a compatible system). We shall rely on the following potential automorphy results when $K$ is totally real.
Theorem 3.4.
[BLGGT14, Theorem C] Let $K$ be a totally real field, $\ell\geq 2(n+1)$ be a rational prime, and
$\rho_{\ell}:\mathrm{Gal}_{K}\to\mathrm{GL}_{n}(\overline{\mathbb{Q}}_{\ell})$
be an $\ell$-adic Galois representation of $K$. Suppose the following conditions are satisfied.
  1. (1)
    (Unramified almost everywhere) $\rho_{\ell}$ is unramified at all but finitely many primes.
  2. (2)
    (Odd essential self-duality) Either $\rho_{\ell}$ maps to $\mathrm{GSp}_{n}$ with totally odd multiplier or it maps to $\mathrm{GO}_{n}$ with totally even multiplier.
  3. (3)
    (Potential diagonalizability and regularity) $\rho_{\ell}$ is potentially diagonalizable (and hence potentially crystalline) at each prime $v$ of $K$ above $\ell$ and for each embedding $\tau:K\to\overline{\mathbb{Q}}_{\ell}$ it has $n$ distinct $\tau$-Hodge-Tate numbers.
  4. (4)
    (Irreducibility) $\overline{\rho}_{\ell}^{\mathrm{ss}}|_{\mathrm{Gal}_{K(\zeta_{\ell})}}$ is irreducible, where $\zeta_{\ell}:=e^{2\pi i/\ell}$ the primitive $\ell$th root of unity.
Then we can find a finite Galois totally real extension $K^{\prime}/K$ such that $\rho_{\ell}|_{\mathrm{Gal}_{K^{\prime}}}$ is attached to a regular algebraic polarized cuspidal automorphic representation of $\mathrm{GL}_{n}(\mathbb{A}_{K^{\prime}})$. Moreover, $\rho_{\ell}$ is part of a strictly compatible system of $K$.
[…]
Proposition 3.5.
[Hui23b, Proposition 2.12(b)] Let $K$ be a totally real field and
$\{\rho_{\lambda}:\mathrm{Gal}_{K}\to\mathrm{GL}_{n}(E_{\lambda})\}_{\lambda}$
be a $E$-rational strictly compatible system. For almost all $\lambda\in\Sigma_{E}$, if $\sigma_{\lambda}$ is a regular three-dimensional subrepresentation of $\rho_{\lambda}\otimes\overline{E}_{\lambda}$ such that the derived subgroup $\mathbf{G}_{\sigma_{\lambda}}^{\mathrm{der}}$ of its algebraic monodromy group is $\mathrm{SO}_{3}$ (as a group embedded in $\mathrm{GL_{3}}$), then there is a finite Galois totally real extension $K^{\prime}/K$ such that $\sigma_{\lambda}|_{\mathrm{Gal}_{K^{\prime}}}$ is attached to a regular algebraic polarized cuspidal automorphic representation of $\mathrm{GL}_{3}(\mathbb{A}_{K^{\prime}})$. Moreover, such $\sigma_{\lambda}$ is a part of a strictly compatible system of $K$.
3.4. Big image results of Galois subrepresentations
We present some big image results for type $A$ Galois subrepresentations in certain ($\mathrm{GL}_{n}(E_{\lambda})$-valued) semisimple compatible system.
Theorem 3.6.
[Hui23a, Theorem 1.2, Theorem 3.12(v)] Let $\{\rho_{\lambda}:\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(E_{\lambda})\}$ be a semisimple compatible system of a number field $K$. Suppose there exist some integers $N_{1},N_{2}\geq 0$ and finite extension $K^{\prime}/K$ such that the following conditions hold.
  1. (a)
    (Bounded tame inertia weights): for almost all $\lambda$ and each finite place $v$ of $K$ above $\ell$, the tame inertia weights of the local representation $(\bar{\rho}_{\lambda}^{\operatorname{ss}}\otimes\bar{\epsilon}_{\ell}^{N_{1}})|_{\operatorname{Gal}_{K_{v}}}$ belong to $[0,N_{2}]$.
  2. (b)
    (Potential semistability): for almost all $\lambda$ and each finite place $w$ of $K^{\prime}$ not above $\ell$, the semisimplification of the local representation $\bar{\rho}_{\lambda}^{\operatorname{ss}}|_{\operatorname{Gal}_{K_{w}^{\prime}}}$ is unramified.
For almost all $\lambda\in\Sigma_{E}$, the following assertions hold.
  1. (i)
    If $\sigma_{\lambda}$ is a type $A$ irreducible subrepresentation of $\rho_{\lambda}\otimes\overline{E}_{\lambda}$, then the residual representation $\bar{\sigma}_{\lambda}^{\operatorname{ss}}$ is also irreducible.
  2. (ii)
    If $\sigma_{\lambda}$ is a type $A$ Lie-irreducible subrepresentation of $\rho_{\lambda}\otimes\overline{E}_{\lambda}$, then the restriction $\bar{\sigma}_{\lambda}^{\operatorname{ss}}|_{\operatorname{Gal}_{K^{ab}}}$ is also irreducible, where $K^{ab}$ is the maximal abelian extension of $K$.
Proposition 3.7.
[Hui23b, Proposition 2.11] Let $\{\rho_{\lambda}:\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(E_{\lambda})\}$ be a $E$-rational strictly compatible system of a number field $K$. Then $\{\rho_{\lambda}\}$ satisfies the conditions 3.6(a),(b) for some integers $N_{1},N_{2}\geq 0$ and finite extension $K^{\prime}/K$.
3.5. $\lambda$-independence of formal bi-character
Let $F$ be an algebraically closed field of characteristic zero and $\mathbf{G}\subset\mathrm{GL}_{n,F}$ a reductive subgroup. Denote by $\mathbf{T}$ a maximal torus of $\mathbf{G}$. Then the intersection $\mathbf{T}^{\operatorname{ss}}:=\mathbf{G}^{\operatorname{der}}\cap\mathbf{T}$ is a maximal torus of the semisimple group $\mathbf{G}^{\operatorname{der}}$. The following definition is independent of the choice of the maximal torus $\mathbf{T}$.
Definition 3.8.
Let $\mathbf{G}\subset\mathrm{GL}_{n,F}$ be a reductive subgroup.
  1. (1)
    The formal character of $\mathbf{G}$ is defined as the conjugacy class of the torus $\mathbf{T}$ in $\mathrm{GL}_{n,F}$.
  2. (2)
    The formal bi-character of $\mathbf{G}$ is defined as the conjugacy class of the chain of subtori $\mathbf{T}^{\operatorname{ss}}\subset\mathbf{T}$ in $\mathrm{GL}_{n,F}$.
Suppose that $\mathbf{G}_{1}$ and $\mathbf{G}_{2}$ are reductive subgroups of respectively $\mathrm{GL}_{n,F_{1}}$ and $\mathrm{GL}_{n,F_{2}}$, where $F_{1}$ and $F_{2}$ are algebraically closed fields of characteristic zero. Embed $F_{1}$ and $F_{2}$ into an algebraically closed field $F$. We say that the formal characters (resp. formal bi-characters) of $\mathbf{G}_{1}$ and $\mathbf{G}_{2}$ are the same if this is true for the base change $\mathbf{G}_{1,F}$ and $\mathbf{G}_{2,F}$ in $\mathrm{GL}_{n,F}$. This definition is independent of the choice of the over-field $F$.
We have the following $\lambda$-independence results about algebraic monodromy groups.
Theorem 3.9.
Let $K$ be a number field and $\{\rho_{\lambda}:\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(\overline{E}_{\lambda})\}$ be a semisimple $E$-rational compatible system with algebraic monodromy groups $\{\mathbf{G}_{\lambda}\}$.
  1. (i)
    [Ser81] The component group $\mathbf{G}_{\lambda}/\mathbf{G}_{\lambda}^{\circ}$ is independent of $\lambda$. In particular, the connectedness of $\mathbf{G}_{\lambda}$ is independent of $\lambda$.
  2. (ii)
    [Ser81] The formal character of $\mathbf{G}_{\lambda}\subset\mathrm{GL}_{n,\overline{E}_{\lambda}}$ is independent of $\lambda$. In particular, the rank of $\mathbf{G}_{\lambda}$ is independent of $\lambda$.
  3. (iii)
    [Hui13, Theorem 3.19, Remark 3.22] The formal bi-character of $\mathbf{G}_{\lambda}\subset\mathrm{GL}_{n,\overline{E}_{\lambda}}$ is independent of $\lambda$. In particular, the formal character of $\mathbf{G}_{\lambda}^{\operatorname{der}}\subset\mathrm{GL}_{n,\overline{E}_{\lambda}}$ (resp. semisimple rank of $\mathbf{G}_{\lambda}$) is independent of $\lambda$.
Remark 3.10.
For a compatible system $\{\rho_{\lambda}\}$ with algebraic monodromy groups $\{\mathbf{G}_{\lambda}\}$, the semisimplification $\{\rho_{\lambda}^{\operatorname{ss}}\}$ is a semisimple compatible system with algebraic monodromy groups $\{\mathbf{G}_{\lambda}/\mathbf{U}_{\lambda}\}$, where $\mathbf{U}_{\lambda}$ is the unipotent radical of $\mathbf{G}_{\lambda}$. Since $\mathbf{U}_{\lambda}$ is always connected, it follows that $\mathbf{G}_{\lambda}/\mathbf{G}_{\lambda}^{\circ}$ is independent of $\lambda$ by Theorem 3.9(i).
A compatible system $\{\rho_{\lambda}\}$ is said to be connected if the algebraic monodromy group $\mathbf{G}_{\lambda}$ is connected for some $\lambda$, equivalently, for all $\lambda$ by Remark 3.10. For any compatible system $\{\rho_{\lambda}\}$ (of $K$), there exists a finite extension $L/K$ such that the restriction $\{\rho_{\lambda}|_{\operatorname{Gal}_{L}}\}$ is connected.
3.6. A refinement of Theorem 3.9(ii),(iii)
[…]
(17) $\{\hat{\rho}_{\lambda}:=\bigoplus_{i=1}^{k}\rho_{\lambda}^{(i)}:\operatorname{Gal}_{K}\to\prod_{i=1}^{k}\mathrm{GL}_{n_{i}}(\overline{E}_{\lambda})\}$
given by direct sum and let
$\{\hat{\mathbf{G}}_{\lambda}\subset\prod_{i=1}^{k}\mathrm{GL}_{n_{i},\overline{E}_{\lambda}}\}$
be the system of algebraic monodromy groups. For all $\lambda\in\Sigma_{E}$, let
$\hat{\mathbf{T}}_{\lambda}^{\operatorname{ss}}\subset\hat{\mathbf{T}}_{\lambda}$
be a chain of subtori in $\prod_{i=1}^{k}\mathrm{GL}_{n_{i},\overline{E}_{\lambda}}$ where $\hat{\mathbf{T}}_{\lambda}$ is a maximal torus of $\hat{\mathbf{G}}_{\lambda}$ and $\hat{\mathbf{T}}_{\lambda}^{\operatorname{ss}}$ is a maximal torus of $\hat{\mathbf{G}}_{\lambda}^{\operatorname{der}}$.
Theorem 3.11.
After the base change $\iota_{\lambda}:\overline{E}_{\lambda}\stackrel{{\scriptstyle\simeq}}{{\rightarrow}}\mathbb{C}$ for all $\lambda\in\Sigma_{E}$,
  1. (i)
    the conjugacy class of subtorus $\hat{\mathbf{T}}_{\lambda,\mathbb{C}}$ in $\prod_{i=1}^{k}\mathrm{GL}_{n_{i},\mathbb{C}}$ is independent of $\lambda$;
  2. (ii)
    the conjugacy class of the chain $\hat{\mathbf{T}}_{\lambda,\mathbb{C}}^{\operatorname{ss}}\subset\hat{\mathbf{T}}_{\lambda,\mathbb{C}}$ in $\prod_{i=1}^{k}\mathrm{GL}_{n_{i},\mathbb{C}}$ is independent of $\lambda$.
Proof.
We may assume that (17) is connected by taking a finite extension of $K$ and the chain $\hat{\mathbf{T}}_{\lambda,\mathbb{C}}^{\operatorname{ss}}\subset\hat{\mathbf{T}}_{\lambda,\mathbb{C}}$ is contained in the diagonal $\prod_{i=1}^{k}\mathbb{G}_{m}^{n_{i}}$ of $\prod_{i=1}^{k}\mathrm{GL}_{n_{i},\mathbb{C}}$ by conjugation.
(i). The idea (as in [Ser81]) is to look at the morphism
$\mathrm{Char}:=\prod_{i=1}^{k}\mathrm{Char}^{(i)}:\prod_{i=1}^{k}\mathrm{GL}_{n_{i},\mathbb{C}}\to\prod_{i=1}^{k}(\mathbb{G}_{a}^{n_{i}-1}\times\mathbb{G}_{m})$
where $\mathrm{Char}^{(i)}(A)=(a_{1},...,a_{n_{i}})$ if $A\in\mathrm{GL}_{n_{i}}(\mathbb{C})$ and $\det(A-T\cdot\mathrm{id})=\sum_{j=0}^{n_{i}}a_{j}T^{n_{i}-j}$. By the compatibility conditions 3.1(a),(b) and Chebotarev’s density theorem, $\mathrm{Char}(\hat{\mathbf{G}}_{\lambda,\mathbb{C}})$ is independent of $\lambda$. Since $\hat{\mathbf{T}}_{\lambda,\mathbb{C}}$ is a maximal torus of the connected $\hat{\mathbf{G}}_{\lambda,\mathbb{C}}$, it follows that $\mathrm{Char}(\hat{\mathbf{T}}_{\lambda,\mathbb{C}})=\mathrm{Char}(\hat{\mathbf{G}}_{\lambda,\mathbb{C}})$ is independent of $\lambda$. Together with the facts that $\hat{\mathbf{T}}_{\lambda,\mathbb{C}}\subset\prod_{i=1}^{k}\mathbb{G}_{m}^{n_{i}}$ (the diagonal) is a closed irreducible subvariety and the restriction $\mathrm{Char}|_{\prod_{i=1}^{k}\mathbb{G}_{m}^{n_{i}}}$ is a finite morphism, we obtain assertion (i).
(ii). This part follows closely the idea behind [Hui13, Theorem 3.19]. Let $\mathfrak{m}$ be a modulus of $K$ and $S_{\mathfrak{m}}$ be the Serre group [^4] The Serre group $S_{\mathfrak{m}}$ is a diagonalizable group defined over $\mathbb{Q}$ and its dimension depends only on $K$, i.e., independent of $\mathfrak{m}$. of $K$ with respect $\mathfrak{m}$. Associated to $S_{\mathfrak{m}}$ a family of abelian $\ell$-adic representations
$\{\alpha_{\ell}:\operatorname{Gal}_{K}\to S_{\mathfrak{m}}(\mathbb{Q}_{\ell})\}_{\ell\in\Sigma_{\mathbb{Q}}}$
such that for any $E$-morphism $\phi:S_{\mathfrak{m},E}\to\mathrm{GL}_{m,E}$, the composition
$\{\phi_{\lambda}:\operatorname{Gal}_{K}\stackrel{{\scriptstyle\alpha_{\ell}}}{{\rightarrow}}S_{\mathfrak{m}}(E_{\lambda})\stackrel{{\scriptstyle\phi\otimes_{E}E_{\lambda}}}{{\longrightarrow}}\mathrm{GL}_{m}(E_{\lambda})\subset\mathrm{GL}_{m}(\overline{E}_{\lambda})\}_{\lambda\in\Sigma_{E}}$
is an abelian semisimple $E$-rational compatible system [Ser98, Chapter II].
Assume $\phi$ is now faithful. Consider the semisimple compatible system $\{\hat{\rho}_{\lambda}\oplus\phi_{\lambda}\}$ (a direct sum of $k+1$ compatible systems) and let
$\{\widetilde{\mathbf{G}}_{\lambda}\subset(\prod_{i=1}^{k}\mathrm{GL}_{n_{i},\overline{E}_{\lambda}})\times\mathrm{GL}_{m,\overline{E}_{\lambda}}\}$
be the system of algebraic monodromy groups. For all $\lambda$, let $\widetilde{\mathbf{T}}_{\lambda}$ be a subtori in $(\prod_{i=1}^{k}\mathrm{GL}_{n_{i},\overline{E}_{\lambda}})\times\mathrm{GL}_{m,\overline{E}_{\lambda}}$ such that $\widetilde{\mathbf{T}}_{\lambda}$ is a maximal torus of $\widetilde{\mathbf{G}}_{\lambda}$. It follows from assertion (i) that
(18) $\text{the conjugacy class of $\widetilde{\mathbf{T}}_{\lambda,\mathbb{C}}$ in $(\prod_{i=1}^{k}\mathrm{GL}_{n_{i},\mathbb{C}})\times\mathrm{GL}_{m,\mathbb{C}}$ is independent of $\lambda$}.$
Let $\pi_{1}$ (resp. $\pi_{2}$) be the projection map of $(\prod_{i=1}^{k}\mathrm{GL}_{n_{i},\mathbb{C}})\times\mathrm{GL}_{m,\mathbb{C}}$ to the first $k$ factors (resp. to the last factor). According to [Hui13, Remark 3.22] and the proof of [Hui13, Theorem 3.19], the chain
(19) $\mathrm{Ker}(\pi_{2}|_{\widetilde{\mathbf{T}}_{\lambda,\mathbb{C}}})^{\circ}\subset\pi_{1}(\widetilde{\mathbf{T}}_{\lambda,\mathbb{C}})$
is the formal bi-character of $\pi_{1}(\widetilde{\mathbf{G}}_{\lambda,\mathbb{C}})=\hat{\mathbf{G}}_{\lambda,\mathbb{C}}$ for all $\lambda$. By (18), the conjugacy class of (19) in $\prod_{i=1}^{k}\mathrm{GL}_{n_{i},\mathbb{C}}$ is independent of $\lambda$. ∎
Suppose $k=2$ in (17) and denote by $\pi_{1}:\mathrm{GL}_{n_{1},\overline{E}_{\lambda}}\times\mathrm{GL}_{n_{2},\overline{E}_{\lambda}}\to\mathrm{GL}_{n_{1},\overline{E}_{\lambda}}$ the projection to the first factor.
Corollary 3.12.
If $\pi_{1}:\hat{\mathbf{G}}_{\lambda_{0}}\to\pi_{1}(\hat{\mathbf{G}}_{\lambda_{0}})$ is an isomorphism for some $\lambda_{0}$, then the surjection
(20) $\pi_{1}:\hat{\mathbf{G}}_{\lambda}\to\pi_{1}(\hat{\mathbf{G}}_{\lambda})$
is an isomorphism for all $\lambda$.
Proof.
First, note that $\pi_{1}(\hat{\mathbf{G}}_{\lambda})$ is the algebraic monodromy group of $\rho_{\lambda}^{(1)}$ for all $\lambda$. Then the condition that (20) is an isomorphism at $\lambda_{0}$ and Theorem 3.9(i),(ii) imply that the ranks (resp. numbers of components) of $\hat{\mathbf{G}}_{\lambda}$ and $\pi_{1}(\hat{\mathbf{G}}_{\lambda})$ are equal for each $\lambda$. Thus, the kernel $\mathbf{C}_{\lambda}$ of the surjection (20) is a finite normal subgroup of $\hat{\mathbf{G}}_{\lambda}^{\circ}$ for each $\lambda$. We obtain that $\mathbf{C}_{\lambda}\subset\hat{\mathbf{T}}_{\lambda}$ (a maximal torus of $\hat{\mathbf{G}}_{\lambda}$) for each $\lambda$. Since the restriction of (20) to $\hat{\mathbf{T}}_{\lambda}$ is injective at $\lambda=\lambda_{0}$, this is also true for all $\lambda$ by Theorem 3.11(i). Therefore, $\mathbf{C}_{\lambda}$ is trivial for all $\lambda$. ∎
3.7. Invariance of roots and $\lambda$-independence
Suppose $\{\rho_{\lambda}:\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(\overline{E}_{\lambda})\}$ is a connected semisimple $E$-rational compatible system with algebraic monodromy groups $\{\mathbf{G}_{\lambda}\}$. After the base change $\iota_{\lambda}:\overline{E}_{\lambda}\stackrel{{\scriptstyle\simeq}}{{\rightarrow}}\mathbb{C}$ for all $\lambda\in\Sigma_{E}$, we obtain a system of complex connected reductive subgroups
$\{\mathbf{G}_{\lambda,\mathbb{C}}\subset\mathrm{GL}_{n,\mathbb{C}}\}.$
Up to $\mathrm{GL}_{n,\mathbb{C}}$-conjugation, there exists a chain of subtori
(21) $\mathbf{T}_{\mathbb{C}}^{\operatorname{ss}}\subset\mathbf{T}_{\mathbb{C}}$
in $\mathrm{GL}_{n,\mathbb{C}}$ such that $\mathbf{T}_{\mathbb{C}}$ is a maximal torus of $\mathbf{G}_{\lambda,\mathbb{C}}$ for all $\lambda$ and $\mathbf{T}_{\mathbb{C}}^{\operatorname{ss}}$ is a maximal torus of $\mathbf{G}_{\lambda,\mathbb{C}}^{\operatorname{der}}$ for all $\lambda$ by Theorem 3.9(iii). We give a crucial criterion, called the invariance of roots, for the connected reductive subgroups $\{\mathbf{G}_{\lambda,\mathbb{C}}\}$ to be conjugate in $\mathrm{GL}_{n,\mathbb{C}}$. Denote by $\mathbb{X}(\mathbf{T}_{\mathbb{C}}^{\operatorname{ss}}):=\mathrm{Hom}(\mathbf{T}_{\mathbb{C}}^{\operatorname{ss}},\mathbb{G}_{m})$ the character group of $\mathbf{T}_{\mathbb{C}}^{\operatorname{ss}}$.
Proposition 3.13.
[Hui18, Corollary 3.9] For two primes $\lambda_{1}$ and $\lambda_{2}\in\Sigma_{E}$, let $\Phi_{1}\subset\mathbb{X}(\mathbf{T}_{\mathbb{C}}^{\operatorname{ss}})$ and $\Phi_{2}\subset\mathbb{X}(\mathbf{T}_{\mathbb{C}}^{\operatorname{ss}})$ be the root systems of the semisimple groups $\mathbf{G}_{\lambda_{1},\mathbb{C}}^{\operatorname{der}}$ and $\mathbf{G}_{\lambda_{2},\mathbb{C}}^{\operatorname{der}}$ with respect to the common maximal torus $\mathbf{T}_{\mathbb{C}}^{\operatorname{ss}}$ in (21). If $\Phi_{1}=\Phi_{2}$, then $\mathbf{G}_{\lambda_{1},\mathbb{C}}$ and $\mathbf{G}_{\lambda_{2},\mathbb{C}}$ are conjugate in $\mathrm{GL}_{n,\mathbb{C}}$.
The main strategy for Theorem 1.7 is to construct some auxiliary compatible system $\{\phi_{\lambda}\}$ so that the invariance of roots criterion is fulfilled.
Proposition 3.14.
Let $\{\rho_{\lambda}:\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(\overline{E}_{\lambda})\}$ be a connected semisimple $E$-rational compatible system with algebraic monodromy groups $\{\mathbf{G}_{\lambda}\subset\mathrm{GL}_{n,\overline{E}_{\lambda}}\}$. Let $\{\phi_{\lambda}:\operatorname{Gal}_{K}\to\mathrm{GL}_{m}(\overline{E}_{\lambda})\}$ be a semisimple $E$-rational compatible system and consider the compatible system $\{\rho_{\lambda}\oplus\phi_{\lambda}\}$ with algebraic monodromy groups $\{\hat{\mathbf{G}}_{\lambda}\subset\mathrm{GL}_{n,\overline{E}_{\lambda}}\times\mathrm{GL}_{m,\overline{E}_{\lambda}}\}$. Let $\pi_{1}$ (resp. $\pi_{2}$) be the projection map of $\mathrm{GL}_{n,\overline{E}_{\lambda}}\times\mathrm{GL}_{m,\overline{E}_{\lambda}}$ to the first (resp. second) factor. Suppose the following conditions hold.
  1. (a)
    For all $\lambda$ (equivalently, for some $\lambda$), the first projection $\pi_{1}:\hat{\mathbf{G}}_{\lambda}\to\mathbf{G}_{\lambda}$ is an isomorphism.
  2. (b)
    For all $\lambda$, the composition $\mathbf{G}_{\lambda}\stackrel{{\scriptstyle\pi_{1}^{-1}}}{{\longrightarrow}}\hat{\mathbf{G}}_{\lambda}\stackrel{{\scriptstyle\pi_{2}}}{{\rightarrow}}\mathrm{GL}_{m,\overline{E}_{\lambda}}$ is the adjoint representation on the semisimple part $\mathrm{Lie}(\mathbf{G}_{\lambda}^{\operatorname{der}})$, where $\pi_{1}^{-1}$ means the inverse of the isomorphism in (a).
After the base change $\overline{E}_{\lambda}\simeq\mathbb{C}$, the conjugacy class of $\mathbf{G}_{\lambda,\mathbb{C}}$ in $\mathrm{GL}_{n,\mathbb{C}}$ is independent of $\lambda$.
Proof.
In the condition (a), Corollary 3.12 implies that the assertion “$\pi_{1}:\hat{\mathbf{G}}_{\lambda}\to\mathbf{G}_{\lambda}$ is an isomorphism” is independent of $\lambda$.
Up to $\mathrm{GL}_{n,\mathbb{C}}\times\mathrm{GL}_{m,\mathbb{C}}$-conjugation, there exists a chain of subtori
$\hat{\mathbf{T}}_{\mathbb{C}}^{\operatorname{ss}}\subset\hat{\mathbf{T}}_{\mathbb{C}}$
in $\mathrm{GL}_{n,\mathbb{C}}\times\mathrm{GL}_{m,\mathbb{C}}$ such that $\hat{\mathbf{T}}_{\mathbb{C}}$ (resp. $\hat{\mathbf{T}}_{\mathbb{C}}^{\operatorname{ss}}$) is a maximal torus of $\hat{\mathbf{G}}_{\lambda,\mathbb{C}}$ (resp. $\hat{\mathbf{G}}_{\lambda,\mathbb{C}}^{\operatorname{der}}$) for all $\lambda$ by Theorem 3.11(ii). Then we may take the chain (21) as
$\pi_{1}(\hat{\mathbf{T}}_{\mathbb{C}}^{\operatorname{ss}})\subset\pi_{1}(\hat{\mathbf{T}}_{\mathbb{C}}).$
By (b), the root system of $\mathbf{G}_{\lambda,\mathbb{C}}^{\operatorname{der}}$ with respect to $\pi_{1}(\hat{\mathbf{T}}_{\mathbb{C}}^{\operatorname{ss}})$ is given by the weights of
$\pi_{1}(\hat{\mathbf{T}}_{\mathbb{C}}^{\operatorname{ss}})\stackrel{{\scriptstyle\pi_{1}^{-1}}}{{\longrightarrow}}\hat{\mathbf{T}}_{\mathbb{C}}^{\operatorname{ss}}\stackrel{{\scriptstyle\pi_{2}}}{{\longrightarrow}}\mathrm{GL}_{m,\mathbb{C}},$
which is independent of $\lambda$. Therefore, the conjugacy class of $\mathbf{G}_{\lambda,\mathbb{C}}$ in $\mathrm{GL}_{n,\mathbb{C}}$ is independent of $\lambda$ by Proposition 3.13. ∎
3.8. Hodge-Tate lift
Let $F$ be a finite field extension of $\mathbb{Q}_{\ell}$ and $\mathbf{H}$ a linear algebraic group defined over $\overline{\mathbb{Q}}_{\ell}$ in this subsection. An $\ell$-adic representation $\rho:\mathrm{Gal}_{F}\to\mathbf{H}(\overline{\mathbb{Q}}_{\ell})$ is said to be Hodge-Tate if for any representation $\mathbf{H}\to\mathrm{GL}_{k,\overline{\mathbb{Q}}_{\ell}}$, the composition
$\mathrm{Gal}_{F}\stackrel{{\scriptstyle\rho}}{{\rightarrow}}\mathbf{H}(\overline{\mathbb{Q}}_{\ell})\to\mathrm{GL}_{k}(\overline{\mathbb{Q}}_{\ell})$
is a Hodge-Tate representation.
Proposition 3.15.
[Pat19, Corollary 3.2.12] Let $\mathbf{H}^{\prime}\rightarrow\mathbf{H}$ be a surjection of $\overline{\mathbb{Q}}_{\ell}$-linear algebraic groups with the central torus kernel and $\rho:\mathrm{Gal}_{F}\to\mathbf{H}(\overline{\mathbb{Q}}_{\ell})$ be a Hodge-Tate representation of $F$. Then there exists a Hodge-Tate lift $\tilde{\rho}:\mathrm{Gal}_{F}\to\mathbf{H}^{\prime}(\overline{\mathbb{Q}}_{\ell})$ so that the following diagram commutes.
${{\mathbf{H}^{\prime}(\overline{\mathbb{Q}}_{\ell})}}$${{\mathrm{Gal}_{F}}}$${{\mathbf{H}(\overline{\mathbb{Q}}_{\ell})}}$$\scriptstyle{\tilde{\rho}}$$\scriptstyle{\rho}$
Proposition 3.16.
Let $\rho_{0}:\mathrm{Gal}_{F}\to\mathrm{GL}_{k}(\overline{\mathbb{Q}}_{\ell})$ be an $\ell$-adic Hodge-Tate representation such that the $\tau$-Hodge-Tate numbers are distinct for any embedding $\tau:F\to\overline{\mathbb{Q}}_{\ell}$. Suppose the following conditions hold.
  1. (a)
    $\rho_{0}$ factors through a connected reductive subgroup $\mathbf{H}\subset\mathrm{GL}_{k,\overline{\mathbb{Q}}_{\ell}}$ of semisimple rank $r_{0}\in\mathbb{N}$.
  2. (b)
    $\mathbf{H}$ is irreducible on the ambient space.
  3. (c)
    $\mathbf{H}$ contains the homothety $\mathbb{G}_{m}$ of $\mathrm{GL}_{k,\overline{\mathbb{Q}}_{\ell}}$.
  4. (d)
    $\mathbf{H}$ is of type $A_{1}$, i.e., $\mathrm{Lie}(\mathbf{H}^{\operatorname{der}})$ has only $A_{1}$ factors.
Then there exist a surjective morphism $\mathbf{H}^{\prime}:=\prod_{j=1}^{r_{0}}\mathrm{GL}_{2,\overline{\mathbb{Q}}_{\ell}}\to\mathbf{H}$ with the central torus kernel and a Hodge-Tate lifting
$\tilde{\rho_{0}}:=\bigoplus_{j=1}^{r_{0}}(f_{j}:\mathrm{Gal}_{F}\to\mathrm{GL}_{2}(\overline{\mathbb{Q}}_{\ell}))$
such that $\tau$-Hodge-Tate numbers of $f_{j}$ are distinct for every $1\leq j\leq r_{0}$.
Proof.
This is essentially the same as the proof of [HL24, Lemma 4.9] (using Proposition 3.15). ∎
3.9. Proof of Theorem 1.7
3.9.1.
Now $K$ is totally real. Since $\rho_{\lambda}$ is $\mathrm{GL}_{n}(E_{\lambda})$-valued, the algebraic monodromy group $\mathbf{G}_{\lambda}$ is defined over $E_{\lambda}$. To prove Theorem  1.7, it suffices (by Proposition 3.14) to construct some compatible system $\{\phi_{\lambda}\}$ (see (24)) so that the conditions 3.14(a),(b) hold (enlarging $E$ if necessary). This will be achieved by the potential automorphy result, Proposition 3.5.
[…]
$\psi_{\lambda}:\operatorname{Gal}_{K}\to\mathrm{GL}_{3r}(E_{\lambda})$
such that $\psi_{\lambda}\otimes\overline{E}_{\lambda}$ is the direct sum of $r$ representations $\psi_{\lambda,i}:\operatorname{Gal}_{K}\to\mathrm{GL}_{3}(\overline{E}_{\lambda})$ for $1\leq i\leq r$, where each $\psi_{\lambda,i}$ corresponds to a simple ($A_{1}$-) factor of $\mathrm{Lie}(\mathbf{G}_{\lambda,\overline{E}_{\lambda}}^{\operatorname{der}})$. Since $\mathbf{G}_{\lambda}$ is connected by the condition 1.7(b), the algebraic monodromy group $\mathbf{G}_{\psi_{\lambda,i}}$ (resp. $\mathbf{G}_{\psi_{\lambda}\otimes\overline{E}_{\lambda}}$) is isomorphic to $\mathrm{SO}_{3}$ (resp. $\mathrm{SO}_{3}^{r}$) for all $i$.
Proposition 3.17.
For each $1\leq i\leq r$, the three dimensional representation $\psi_{\lambda,i}$ is regular.
Proof.
Since $\{\rho_{\lambda}\}$ is strictly compatible by the condition 1.7(a), the system $\{\rho_{\lambda}\otimes\rho_{\lambda}^{\vee}\}$ is also strictly compatible. As $\psi_{\lambda}$ is a subrepresentation of $\rho_{\lambda}\otimes\rho_{\lambda}^{\vee}$, it is de Rham at places $v$ above $\ell$. It remains to show that $\psi_{\lambda,i}|_{\operatorname{Gal}_{K_{v}}}$ has distinct $\tau$-Hodge-Tate numbers for any embedding $\tau:K_{v}\to\overline{E}_{\lambda}$ (see Definition 3.3).
Take $F:=K_{v}$ and $\overline{E}_{\lambda}\simeq\overline{\mathbb{Q}}_{\ell}$ a field isomorphism. For each non-abelian irreducible subrepresentation $\rho_{0}$ of $\rho_{\lambda}\otimes\overline{E}_{\lambda}$, the restriction
(22) $\rho_{0}:\operatorname{Gal}_{F}\to\mathrm{GL}_{k}(\overline{\mathbb{Q}}_{\ell})$
has distinct $\tau$-Hodge-Tate numbers by the condition 1.7(a). Moreover, (22) satisfies the conditions 3.16(a)–(d) with $\mathbf{H}:=\mathbf{G}_{\rho_{0}}\mathbb{G}_{m}$, where $\mathbb{G}_{m}$ denotes the homothety of $\mathrm{GL}_{k,\overline{\mathbb{Q}}_{\ell}}$. Hence, we obtain $f_{1},...,f_{r_{0}}$ by Proposition 3.16, where $r_{0}\in\mathbb{N}$ is the semisimple rank of $\mathbf{G}_{\rho_{0}}$.
Since $\psi_{\lambda,i}$ is a simple $A_{1}$-factor of $\psi_{\lambda}\otimes\overline{\mathbb{Q}}_{\ell}=\mathrm{Lie}(\mathbf{G}_{\lambda,\overline{\mathbb{Q}}_{\ell}}^{\operatorname{der}})$, there exists an irreducible $\rho_{0}$ (as above) such that $\psi_{\lambda,i}$ is injective via the natural $\operatorname{Gal}_{F}$-equivariant morphism
$\mathrm{Lie}(\mathbf{G}_{\lambda,\overline{\mathbb{Q}}_{\ell}}^{\operatorname{der}})\to\mathrm{Lie}(\mathbf{G}_{\rho_{0}}^{\operatorname{der}})=\mathrm{Lie}(\mathbf{H}^{\operatorname{der}})\simeq\mathrm{Lie}(\mathbf{H}^{\prime\operatorname{der}}),$
where the first arrow comes from $\rho_{0}$-projection and the last isomorphism holds because the kernel of $\mathbf{H}^{\prime}\twoheadrightarrow\mathbf{H}$ is a central torus by Proposition 3.16. Hence, we find some $f_{j}:\operatorname{Gal}_{F}\to\mathrm{GL}_{2}(\overline{\mathbb{Q}}_{\ell})$ such that the restriction
(23) $\psi_{\lambda,i}|_{\operatorname{Gal}_{F}}\simeq\mathrm{ad}^{0}(f_{j})$
as $\operatorname{Gal}_{F}$-representations, where $\mathrm{ad}^{0}(f_{j})$ denotes the trace zero part of $f_{j}\otimes f_{j}^{\vee}=\operatorname{End}(f_{j})$. Since the $\tau$-Hodge-Tate numbers of $f_{j}$ are distinct by Proposition 3.16, the same holds for $\psi_{\lambda,i}|_{\operatorname{Gal}_{F}}$ by (23). ∎
3.9.2.
[…]
(24) $\{\phi_{\lambda}:=\bigoplus_{i=1}^{r}\phi_{i,\lambda}:\operatorname{Gal}_{K}\to\prod_{i=1}^{r}\mathrm{GL}_{3}(\overline{E}_{\lambda})\}_{\lambda\in\Sigma_{E}}$
of the totally real $K$ such that $\phi_{\lambda_{0}}=\psi_{\lambda_{0}}$.
Proposition 3.18.
For all $\lambda$, the algebraic monodromy group $\mathbf{G}_{\phi_{\lambda}}$ of $\phi_{\lambda}$ is $\mathrm{SO}_{3}^{r}\subset\mathrm{GL}_{3}^{r}$.
Proof.
Note that the assertion holds at $\lambda_{0}$. In particular, the algebraic monodromy group of $\phi_{i,\lambda_{0}}$ satisfies
(25) $\mathbf{G}_{\phi_{i,\lambda_{0}}}=\mathrm{SO}_{3}\subset\mathrm{GL}_{3}$
for every $i$. As Proposition 3.5 asserts that the restriction $\{\phi_{i,\lambda}|_{\operatorname{Gal}_{K_{i}}}\}$ is automorphic for some totally real extension $K_{i}/K$, the algebraic monodromy group of $\phi_{i,\lambda}|_{\operatorname{Gal}_{K_{i}}}$ is independent of $\lambda$ (see [Hui23a, Proposition 4.16(iii)]), which must be $\mathrm{SO}_{3}\subset\mathrm{GL}_{3}$ by (25). By (25) again and Theorem 3.9, we obtain
$\mathbf{G}_{\phi_{i,\lambda}}=\mathrm{SO}_{3}\subset\mathrm{GL}_{3}$
for all $\lambda$ and $i$, which implies that
(26) $\mathbf{G}_{\phi_{\lambda}}\subset\mathrm{SO}_{3}^{r}$
in $\mathrm{GL}_{3}^{r}$ for all $\lambda$. Since (26) is an equality at $\lambda_{0}$, it follows from Theorem 3.9 that (26) is an equality for all $\lambda$. We are done. ∎
3.9.3.
With $\{\phi_{\lambda}\}$ at hand, we are in the setting of Proposition 3.14. To complete the proof of Theorem 1.7, it remains to verify the conditions 3.14(a),(b). The algebraic monodromy group $\hat{\mathbf{G}}_{\lambda}$ of $\rho_{\lambda}\oplus\phi_{\lambda}$ is contained in the product $\mathbf{G}_{\lambda}\times\mathbf{G}_{\phi_{\lambda}}\subset\mathrm{GL}_{n}\times\mathrm{GL}_{3r}$. Let $\pi_{1}$ and $\pi_{2}$ be respectively, the projection to the first and second factor. Since $\phi_{\lambda_{0}}$ is given by the adjoint action on $\mathrm{Lie}(\mathbf{G}_{\lambda_{0}}^{\operatorname{der}})$, the condition 3.14(a) holds (at $\lambda=\lambda_{0}$). Since the image of
(27) $\mathbf{G}_{\lambda}\stackrel{{\scriptstyle\pi_{1}^{-1}}}{{\longrightarrow}}\hat{\mathbf{G}}_{\lambda}\stackrel{{\scriptstyle\pi_{2}}}{{\rightarrow}}\mathrm{GL}_{3r,\overline{E}_{\lambda}}$
is $\mathrm{SO}_{3}^{r}\subset\mathrm{GL}_{3}^{r}$ (Proposition 3.18) and $r$ is the semisimple rank of $\mathbf{G}_{\lambda}$, (27) can only be the adjoint representation on the semisimple part $\mathrm{Lie}(\mathbf{G}_{\lambda}^{\operatorname{der}})$, i.e., the condition 3.14(b) holds for all $\lambda$.
3.10. Proof of Corollary 1.9
The compatible system $\{\rho_{\lambda}\}$ is rectangular by the condition 1.9(c). Hence, Corollary 1.9(i) follows immediately from Theorem 1.7. Since $\rho_{\lambda_{0}}$ is absolutely irreducible (1.9(c)), Corollary 1.9(ii) follows directly from Corollary 1.9(i). Since $\mathbf{G}_{\lambda,\mathbb{C}}$ is of type $A_{1}$ for all $\lambda$, (Corollary 1.9(i) and 1.9(c)), Corollary 1.9(iii) follows directly from Corollary 1.9(ii), Proposition 3.7, and Theorem 3.6(i).
Acknowledgments
C.Y. Hui would like to thank Kei Yuen Chan, Zachary Feng, Dmitri Whitmore, Kayue Daniel Wong, and Jun Yu for their interests and feedback in the article. C.Y. Hui was partially supported by Hong Kong RGC (no. 17314522), NSFC (no. 12222120), and a Humboldt Research Fellowship.